### v1

# Half canonical rings of Gorenstein spin curves of genus two

Abstract.
We introduce the notion of generalised Gorenstein spin structure on a curve and we give an explicit description of the associated section ring for curves of genus two with ample canonical bundle, obtaining five different formats.
Key words and phrases:
spin structure, Gorenstein curve, section ring
1991 Mathematics Subject Classification:
14H42; 14H45, 14H20
1. Introduction
Recall that a spin structure (or theta-characteristic) on a smooth curve $C$ is a line bundle ${\mathcal{L}}$ such that there is an isomorphism $\mu\colon{\mathcal{L}}\otimes{\mathcal{L}}\to\omega_{C}$. The interest in such structures originated in the classical study of theta functions in the 19th century and they were taken up again in 1970’s in the context of the study of moduli spaces of curves (see [Far12]). They also occur in the study of algebraic surfaces: if $X$ is a smooth projective surface and $C\in|K_{X}|$ is a smooth curve in the canonical linear system, then ${\mathcal{L}}=K_{X}|_{C}$ is a theta-characteristic by the adjunction formula.
Both in applications to moduli and to algebraic surfaces it soon became necessary to consider degenerations of the above description, with both the curve and the line bundle possibly acquiring singularities. In the context of moduli of curves this problem has seen some attention [Cor89, Jar98, Lud10]. However, the application to algebraic surfaces, that sparked our interest in these structures and will be explained in a subsequent paper, requires a different approach than what was available in the literature.
More precisely, the curves $C$ we are led to consider are Gorenstein curves $C$ of genus two with ample canonical bundle $\omega_{C}$. Such curves, if reduced, can be classified into two types, Type A and Type B, according to the image of the bicanonical map (Proposition 3.3).
A generalised Gorenstein spin (ggs) structure on $C$ is given by a torsion-free sheaf ${\mathcal{L}}$ everywhere of rank one such that $\chi({\mathcal{L}})=0$ and a multiplication map $\mu\colon{\mathcal{L}}\operatorname{\otimes}{\mathcal{L}}\to\omega_{C}$ which is an isomorphism at the generic points. If ${\mathcal{L}}$ is in fact a line bundle, what we call a regular ggs structure, then this is the direct generalisation of the original notion. We will see in Section 4 that otherwise the ggs structures on $C$ are governed by regular ggs structures on partial normalisations of $C$ in many cases. This leads to a pretty complete understanding of the possible ggs structures on the reduced curves of interest. The non-reduced cases, which presents more complications, is not addressed here.
For stable curves our definition of spin structure coincides with the definition given in [Jar98] but, in view of the application to stable surfaces, we need to include both curves with cusps and curves with elliptic tails in our analysis. As explained in the comments after Definition 1.2 in [FS13], these two classes of curves cannot occur at the same time in a separated moduli problem for curves, so it is not clear how to deal with families in our setup. We have chosen not to pursue this question here.
We then go on and analyse the half-canonical ring of such a ggs structure, that is,
$R(C,\{{\mathcal{L}},\omega_{C}\})=\bigoplus_{n}R_{n}$
where $R_{2n}=H^{0}(C,nK_{C})$ and $R_{2n+1}=H^{0}(C,{\mathcal{L}}(nK_{C}))$. Note that this is not a standard section ring, because the multiplication of two elements of odd degree has to be defined via the multiplication map $\mu$. Leaving the precise statements to Section 5 we prove the following.
Theorem 5.2 (simplified version) —
Let $({\mathcal{L}},\mu)$ be a ggs structure on a reduced Gorenstein curve of genus two with ample canonical bundle. Then the following cases occur:
$h^{0}({\mathcal{L}})$ 0 1 2
Type of ring $A(0)$, $B(0)$ $A(1)$, $B(1)$ $B(2)$
In cases $A(0)$, $A(1)$, and $B(2)$ we get complete intersections, in cases $B(0)$ and $B(1)$ the rings admit a more complicated presentation in codimension six and four.
Note that in each case we can reinterpret the ring as giving a natural embedding of $C$ into a weighted projective space ${\mathbb{P}}$, such that $\omega_{C}={\mathcal{O}}_{\mathbb{P}}(2)|_{C}$ and ${\mathcal{L}}={\mathcal{O}}_{\mathbb{P}}(1)|_{C}/\text{torsion}$.
We hope these considerations to be of independent interest, a direct application to the moduli space of stable surfaces will be given in a subsequent paper.
Acknowledgements
We thank Igor Burban for some help with torsion-free sheaves on curve singularities. S.R. is grateful for support by the DFG. M.F. and R.P. are partially supported by the project PRIN 2022BTA242 “Geometry of algebraic structures: Moduli, Invariants, Deformations” of Italian MUR and members of GNSAGA of INDAM.
2. Notation and preliminary results
2.A. Set-up
We work with schemes of finite type over the complex numbers.
Given a sheaf ${\mathcal{F}}$ on a scheme $X$, we denote by ${\mathcal{F}}^{\vee}=\curly{H}om({\mathcal{F}},\mathcal{O}_{X})$ its dual and by ${\mathcal{F}}^{[m]}$ the $m$-th reflexive power, i.e., the double dual of ${\mathcal{F}}^{\operatorname{\otimes}m}$. The sheaf ${\mathcal{F}}$ is reflexive if ${\mathcal{F}}\cong{\mathcal{F}}^{\vee\vee}$.
A curve $C$ is a Cohen-Macaulay projective scheme of pure dimension 1. It has a dualising sheaf $\omega_{C}$ and we denote by $p_{a}(C)$ the arithmetic genus of $C$, that is, $p_{a}(C)=1-\chi(\mathcal{O}_{C})$. The curve $C$ is Gorenstein if $\omega_{C}$ is invertible; if this is the case $K_{C}$ denotes a canonical divisor such that $\mathcal{O}_{C}(K_{C})\cong\omega_{C}$.
We say that $p\in C$ is a singular point of type $A_{m}$ if the completion of the local ring $\mathcal{O}_{C,p}$ is isomorphic to the ring ${\mathbb{C}}\llbracket x,y\rrbracket/(y^{2}-x^{m+1})$.
A variety is a reduced (possibly reducible) Cohen–Macaulay scheme.
2.B. Rank one torsion-free sheaves on Gorenstein curves
A torsion-free sheaf on a curve $C$ is a coherent sheaf ${\mathcal{F}}$ such that $10pt({\mathcal{F}}_{x})=$ 1 for every closed point $x\in C$; equivalently ${\mathcal{F}}$ has no subsheaves supported on a finite set of points, i.e., ${\mathcal{F}}$ satisfies Serre condition $S_{1}$.
Convention. By a rank one torsion free sheaf on a curve $C$ we mean a torsion free sheaf that has rank one at the general point of any component of $C$.
The degree of a rank one torsion-free sheaf ${\mathcal{F}}$ on a curve $C$ is defined by the formula $\deg({\mathcal{F}})=\chi({\mathcal{F}})-\chi(\mathcal{O})$, so that the Riemann–Roch formula is true by definition.
Remark 2.1
Note that by Serre duality on a Cohen–Macaulay curve $C$ we still have $\chi(\omega_{C})=-\chi({\mathcal{O}}_{C})$, so $\deg\omega_{C}=2p_{a}(C)-2$. If ${\mathcal{F}}\hookrightarrow{\mathcal{F}}^{\prime}$ is an injection with finite cokernel, then $\deg{\mathcal{F}}\leq\deg{\mathcal{F}}^{\prime}$ by the additivity of the Euler characteristic.
In general, every reflexive sheaf on a curve is torsion-free and by a result of Hartshorne these notions coincide for rank one sheaves on a Gorenstein curve.
Lemma 2.2 ([Har94], Prop. 1.6) —
Let $C$ be a Gorenstein curve and ${\mathcal{F}}$ a rank one coherent sheaf on $C$. Then ${\mathcal{F}}$ is reflexive if and only if it is torsion-free.
In particular, for any such sheaf ${\mathcal{F}}$ we have
[…]
Moreover, in this case, if ${\mathcal{F}}$ is a rank one torsion-free sheaf and ${\mathcal{H}}$ is an invertible sheaf then $\deg({\mathcal{F}}\otimes{\mathcal{H}})=\deg({\mathcal{F}})+\deg({\mathcal{H}})$ (cf. [Har94, Prop. 2.8]). We point out a useful Lemma, which relates the behaviour of a rank one torsion free sheaf and its pull-back under a partial normalisation.
Lemma 2.3 —
Let ${\mathcal{F}}$ be a rank one torsion free sheaf on a reduced Gorenstein curve. Then:
  1. (i)
    there are a (uniquely determined) partial normalisation $f\colon\widetilde{C}\to C$ and a line bundle $\widetilde{\mathcal{F}}$ on $\widetilde{C}$ such that ${\mathcal{F}}=f_{*}\widetilde{\mathcal{F}}$;
  2. (ii)
    $\operatorname{Aut}({\mathcal{F}})=\operatorname{Aut}(\widetilde{\mathcal{F}})$. In particular, if $\widetilde{C}$ is connected then $\operatorname{Aut}({\mathcal{F}})={\mathbb{C}}^{*}$.
Proof.
$(i)$ Since at a smooth point of a curve every torsion-free sheaf is free, the question is local at the singular points of $C$. Then the existence of the partial normalisation and the line bundle $\widetilde{\mathcal{F}}$ follows from [Bas63, Prop. 7.2]. For the unicity, note that the natural map $f^{*}{\mathcal{F}}=f^{*}f_{*}\widetilde{\mathcal{F}}\to\widetilde{\mathcal{F}}$ is surjective and therefore it induces an isomorphism $(f^{*}{\mathcal{F}})^{[1]}\cong\widetilde{\mathcal{F}}$, since both sheaves have rank one and $\widetilde{\mathcal{F}}$ is locally free. Hence $\widetilde{\mathcal{F}}$ can be recovered from ${\mathcal{F}}$.
$(ii)$ Automorphisms of ${\mathcal{F}}$, $\widetilde{\mathcal{F}}$ are in bijection. In fact an automorphism of $\widetilde{\mathcal{F}}$ obviously induces an automorphism of ${\mathcal{F}}:=f_{*}\widetilde{\mathcal{F}}$. Conversely an automorphism of ${\mathcal{F}}$ induces an automorphism of $(f^{*}{\mathcal{F}})^{[1]}\cong\widetilde{\mathcal{F}}$. $\square$
The result that follows is certainly well known to experts, but we give a proof for lack of a suitable reference.
Lemma 2.4 —
Let $C$ be a reduced Gorenstein curve and let $A,B\subset C$ be subcurves without common components such that $C=A\cup B$. Then there is an exact sequence:
$0\to\omega_{B}\to\omega_{C}\to\omega_{C}{\raisebox{-0.1pt}{$|_{A}$}}\to 0$
Proof.
Let ${\mathcal{I}}_{A},{\mathcal{I}}_{B}\subset{\mathcal{O}}_{C}$ be the ideals of $A$, $B$, respectively. Both ideals are radical and they satisfy ${\mathcal{I}}_{A}{\mathcal{I}}_{B}=0$. More precisely, ${\mathcal{I}}_{A}$ is the annihilator of ${\mathcal{I}}_{B}$, and viceversa. By relative duality for finite morphisms [Sta23, Section 0AU3] we have an inclusion $\omega_{B}=\curly{H}om_{\mathcal{O}_{C}}({\mathcal{O}}_{B},\omega_{C})\hookrightarrow\omega_{C}$, so it is enough to give an isomorphism ${\mathcal{I}}_{A}\omega_{C}\to\curly{H}om_{\mathcal{O}_{C}}({\mathcal{O}}_{B},\omega_{C})$. In fact, there is a natural map $F\colon{\mathcal{I}}_{A}\omega_{C}\to\curly{H}om_{\mathcal{O}_{C}}({\mathcal{O}}_{B},\omega_{C})$, that maps $\sigma\in{\mathcal{I}}_{A}\omega_{C}$ to the homomorphism that sends $1\in\mathcal{O}_{B}$ to $\sigma$ and, using the fact that $\omega_{C}$ is invertible and ${\mathcal{I}}_{A}$ is the annihilator of ${\mathcal{I}}_{B}$, it is immediate to check that $F$ is an isomorphism. $\square$
For later reference, we recall two results related to sections and cohomology of torsion-free sheaves.
Lemma 2.5 —
Let $C$ be a reduced Gorenstein curve and ${\mathcal{F}}$ a rank one torsion-free sheaf on $C$. If $\deg\left({\mathcal{F}}{\raisebox{-0.1pt}{$|_{B}$}}^{[{1}]}\right)\geq 2p_{a}(B)-1$ for every subcurve $B\subseteq C$, then $H^{1}(C,{\mathcal{F}})=0$.
Proof.
The proof follows by the arguments used in [CFHR99, Thm.1.1].
By Serre duality $H^{1}(C,{\mathcal{F}})^{\vee}\cong\operatorname{Hom}({\mathcal{F}},\omega_{C})$. Assume it is not zero and pick any nonzero map $\varphi\colon{\mathcal{F}}\to\omega_{C}$. By [CFHR99, Lemma 2.4], $\varphi$ comes from a generically surjective map ${\mathcal{F}}{\raisebox{-0.1pt}{$|_{B}$}}\to\omega_{B}$ for a subcurve $B\subseteq C$ and yields an injection ${\mathcal{F}}{\raisebox{-0.1pt}{$|_{B}$}}^{[{1}]}\hookrightarrow\omega_{B}$ whose cokernel has finite length. Therefore $\chi\left({\mathcal{F}}{\raisebox{-0.1pt}{$|_{B}$}}^{[{1}]}\right)\leq\chi(\omega_{B})$, and Remark 2.1 gives $\deg\left({\mathcal{F}}{\raisebox{-0.1pt}{$|_{B}$}}^{[{1}]}\right)\leq 2p_{a}(B)-2$, against the assumptions. $\square$
To conclude this section we recall a slight generalisation of the Base-Point Free Pencil Trick, a way to study the behavior of multiplication of sections on a connected curve (see [Mum70, §1 Thm. 2], [Fra04, Prop. 1.5]).
Proposition 2.6 —
Let $C$ be a connected reduced Gorenstein curve, let ${\mathcal{H}}$ be an invertible sheaf on $C$ such that $|{\mathcal{H}}|$ is a base point free system, and let ${\mathcal{F}}$ be a rank one torsion-free sheaf on $C$. Then the multiplication map
$\rho_{{\mathcal{F}},{\mathcal{H}}}:H^{0}(C,{\mathcal{H}})\otimes H^{0}(C,{\mathcal{F}})\rightarrow H^{0}(C,{\mathcal{H}}\otimes{\mathcal{F}})$
is surjective if one of the following conditions holds
  1. (i)
    $H^{1}(C,{\mathcal{F}}\otimes{\mathcal{H}}^{-1})=0$; by Lemma 2.5 this holds if $\deg\left(({\mathcal{F}}\otimes{\mathcal{H}}^{-1})_{|B}^{[{1}]}\right)\geq 2p_{a}(B)-1$ for every subcurve $B\subseteq C$.
  2. (ii)
    $h^{0}(C,{\mathcal{H}})\geq 3$ and ${\mathcal{F}}\otimes{\mathcal{H}}^{-1}\cong\omega_{C}$.
Moreover, if $h^{0}(C,{\mathcal{H}})=2$ then $\ker\rho_{{\mathcal{F}},{\mathcal{H}}}\cong H^{0}(C,{\mathcal{F}}\otimes{\mathcal{H}}^{-1}).$
Proof.
(i) follows by adapting to our case the arguments used in [Mum70, §1 Thm. 2] (which is stated for any coherent sheaf ${\mathcal{F}}$ on a variety $X$). Indeed, since $|{\mathcal{H}}|$ is a base point free system, we can take a section which yields an inclusion ${\mathcal{F}}\otimes{\mathcal{H}}^{-1}\to{\mathcal{F}}$; denoting by ${\mathcal{G}}$ its cokernel, by our assumptions $\dim(\operatorname{Supp}({\mathcal{G}}))=0$ and we have an exact sequence
$0\to H^{0}(C,{\mathcal{F}}\otimes{\mathcal{H}}^{-1})\otimes H^{0}(C,{\mathcal{H}})\to H^{0}(C,{\mathcal{F}})\otimes H^{0}(C,{\mathcal{H}})\to H^{0}(C,{\mathcal{G}})\otimes H^{0}(C,{\mathcal{H}})\to 0;$
[…]
$(ii)$ follows from [Fra04, Prop. 1.5], since in this case ${\mathcal{F}}$ is invertible.
The last statement is the classical Castelnuovo’s Base-Point Free Pencil Trick as shown in [ACGH85, p. 126]. $\square$
3. Reduced Gorenstein curves of genus two with ample canonical bundle
The contribution of a singular point of a curve to the global genus, compared to the normalisation, is measured by the so-called $\delta$-invariant of the point. Singularities with small value of this invariant have been classified by Stevens [Ste96] and later independently in [Smy11, Appendix A] and [Bat22, Sect. 2]. However, without further conditions, there are infinitely many Gorenstein curves of any fixed genus.
Example 3.1 —
If $C$ is any nodal curve such that every component is isomorphic to ${\mathbb{P}}^{1}$ and the dual graph is a tree, then $C$ is Gorenstein of arithmetic genus zero.
Clearly, every irreducible curve of genus zero is isomorphic to ${\mathbb{P}}^{1}$.
Example 3.2 —
There are three different types of irreducible Gorenstein curves $C$ of arithmetic genus $1$, namely a smooth elliptic curve, or a rational curve with a node $A_{1}$ or a cusp $A_{2}$.
This can either be deduced from Stevens’ classification or by showing that a line bundle of degree three embeds $C$ as a plane cubic reasoning as in Proposition 3.3 below.
There are several ways to understand and describe the curves we are interested in. For sake of brevity, we take a quick route via the canonical ring, where we double the degrees in consistency with later considerations.
Proposition 3.3 —
Let $C$ be a reduced connected Gorenstein curve of arithmetic genus two with ample canonical bundle. Then the canonical ring is isomorphic to one of the following:
Type A:
$R(C,\omega_{C})={\mathbb{C}}[y_{1},y_{2},u]/(u^{2}-f_{12}(y_{1},y_{2})),$ where $\deg(y_{1},y_{2},u)=(2,2,6)$ and $f_{12}$ is a non-zero polynomial of weighted degree $12$.
Type B:
$R(C,\omega_{C})\cong{\mathbb{C}}[y_{1},y_{2},v,u]/(y_{1}y_{2},u^{2}-f_{12}(y_{1},y_{2},v)),$ where $\deg(y_{1},y_{2},v,u)=(2,2,4,6)$ and $f_{12}$ is a polynomial of degree $12$ containing the monomial $v^{3}$ with non-zero coefficient.
Proof.
Note that for a Gorenstein curve the Hilbert series of the canonical ring is determined by the genus and the Riemann–Roch formula, so in our case the series is $H(t)=1+2t^{2}+\sum_{n\geq 2}(2n-1)t^{2n}$. Since
$t^{2}H(t)=t^{2}+2t^{4}+\sum_{n\geq 3}(2n-3)t^{2n},$
we may subtract the latter series from the former to get
$(1-t^{2})H(t)=1+t^{2}+t^{4}+\sum_{n\geq 3}2t^{2n}.$
Repeating the above procedure, we have
$t^{2}(1-t^{2})H(t)=t^{2}+t^{4}+t^{6}+\sum_{n\geq 4}2t^{2n}$
and subtracting yields $(1-t^{2})^{2}H(t)=1+t^{6}$. Hence
$H(t)=\frac{1+t^{6}}{(1-t^{2})^{2}}=\frac{(1-t^{12})}{(1-t^{2})^{2}(1-t^{6})}.$
If $C$ is irreducible, we can proceed as for smooth curves (compare [FPR17, Section 3.1]) and get a canonical ring of type A.
Now assume $C$ is reducible. Note that an ample line bundle has positive degree on any component, so there are at most two components and we can write $C=C_{1}\cup C_{2}$ with $\deg(\omega_{C}|_{C_{i}})=1$.
By Lemma 2.4 there is an exact sequence
(3.4) $0\to\omega_{C_{1}}\to\omega_{C}\to\omega_{C}|_{C_{2}}\to 0$
This implies that $\omega_{C_{i}}\hookrightarrow\omega_{C}|_{C_{i}}$, so by Remark 2.1 we get $\deg\omega_{C_{i}}\leq 0$ by parity and thus $C_{i}$ has arithmetic genus at most one. Therefore, each component is Gorenstein by [Cat82, Lemma 1.19].
Taking global sections in (3.4) after twisting by $(m-1)K_{C}$ with $m\geq 1$, we get an exact sequence
(3.5) $0\to H^{0}(K_{C_{1}}+(m-1)K_{C}|_{C_{1}})\to H^{0}(mK_{C})\to H^{0}(mK_{C}|_{C_{2}})\to 0$
because by Serre duality $H^{1}(K_{C_{1}}+(m-1)K_{C}|_{C_{1}})\to H^{1}(mK_{C})$ is an isomorphism for $m=1$ and both are zero for $m\geq 2$.
Now assume $p_{a}(C_{2})=0$, that is, $C_{2}$ is a smooth rational curve. Then $\omega_{C}|_{C_{2}}={\mathcal{O}}_{C_{2}}(1)$ has two sections, so sequence (3.5) for $m=1$ shows that $h^{0}(K_{C_{1}})=0$, and therefore both components are rational. Also, since $K_{C}|_{C_{i}}$ has degree one, it is base-point free on the components and thus globally base point free, because all sections lift to $C$. By Proposition 2.6 $(i)$ the multiplication map
$H^{0}(C,K_{C})\otimes H^{0}(C,nK_{C})\to H^{0}(C,(n+1)K_{C})$
is onto for $n\geq 3$. For $n=1$ one can check easily that $\langle y_{1}^{2},y_{2}^{2},y_{1}y_{2}\rangle$ span $H^{0}(C,2K_{C})$, where $y_{1},y_{2}$ is a basis of $H^{0}(K_{C})$. For $n=2$ the kernel of the multiplication map has dimension two (generated by the trivial relations $y_{1}\operatorname{\otimes}y_{1}y_{2}-y_{2}\operatorname{\otimes}y_{1}^{2},y_{2}\operatorname{\otimes}y_{1}y_{2}-y_{1}\operatorname{\otimes}y_{2}^{2}$), hence we need a new generator $u\in H^{0}(3K_{C})$.
Concerning the relations, we obtain a non-trivial relation between monomials in our generators in $H^{0}(6K_{C})$, which we can write in the given form by completing the square. Comparing the Hilbert series, we see that we have already found the full ring of type A.
Now we turn to the case where both components are of arithmetic genus one. By [Cat82, Lemma 1.12] the ideal ${\mathcal{O}}_{C_{2}}(-C_{1})$ is invertible and defines a subscheme of length $\deg\omega_{C}|_{C_{2}}-\deg\omega_{C_{2}}=1$, i.e., a single point $p\in C$, which by [Cat82, Prop. 1.10] is a node of $C$. Then sequence (3.5) for $m=1$ shows that $p$ is a base point of the canonical system. Let $y_{i}$ be a generator of the kernel of $H^{0}(K_{C})\to H^{0}(K_{C}|_{C_{i}})$. Then $H^{0}(K_{C})=\langle y_{1},y_{2}\rangle$ such that $y_{i}|_{C_{i}}=0$. In particular, $y_{1}y_{2}=0$ and we can write $H^{0}(2K_{C})=\langle y_{1}^{2},y_{2}^{2},v\rangle$. Note by (3.5), this time for $m=2$, that $y_{1}^{2},v$ restricts to $C_{2}$ as a complete linear system of degree two, thus without base points. Hence also $|2K_{C}|$ has no base points, $v$ is non-zero on both components, and by Proposition 2.6 $(ii)$ for $n=3$ and $(i)$ for $n\geq 4$, we obtain surjections
[…]
for every $n\geq 3$.
Now consider the multiplication map $H^{0}(K_{C})\otimes H^{0}(2K_{C})\to H^{0}(3K_{C})$. The image of this map is generated by the monomials $\{y_{1}^{3},y_{2}^{3},y_{1}v,y_{2}v\}$. We have no linear relations among them since a combination
$a_{1}y_{1}^{3}+a_{2}y_{2}^{3}+b_{1}y_{1}v+b_{2}y_{2}v=0$
when restricted to $C_{1}$ yields $a_{2}y_{2}^{3}+b_{2}y_{2}v=0$, whence $a_{2}=b_{2}=0$, and then restricted to $C_{2}$ gives $a_{1}=b_{1}=0$. Therefore, since $h^{0}(C,3K_{C})=5$ we need a new generator $u$ of degree 6.
Similarly we can show that $H^{0}(4K_{C})$ is generated by the monomials
$\{y_{1}^{4},y_{2}^{4},y_{1}^{2}v,y_{2}^{2}v,v^{2},y_{1}u,y_{2}u\}.$
Summing up, the canonical ring $R(C,K_{C})$ is generated by $\{y_{1},y_{2},v,u\}$.
Concerning the relations, we have trivial relations induced by $y_{1}y_{2}=0$ in degree $\leq 10$, and a new relation in degree $12$, which after completing the square can be written as $u^{2}-f_{12}(y_{1},y_{2},v)$. Note that $f_{12}$ must contain the monomial $v^{3}$ with non-zero coefficient, since otherwise the intersection point $C_{1}\cap C_{2}$ corresponding to the ideal $(y_{1},y_{2},u^{2}-f_{12}(y_{1},y_{2},v))$ would be the point $(0:0:1:0)$ and therefore singular for both $C_{1}$ and $C_{2}$. Comparing the Hilbert series, $\frac{(1-t^{12})(1-t^{4})}{(1-t^{2})^{2}(1-t^{6})(1-t^{4})}=\frac{(1-t^{12})}{(1-t^{2})^{2}(1-t^{6})}$ we see that we have already found the full ring of type B. $\square$
Corollary 3.6 —
Let $C$ be a reduced connected Gorenstein curve of arithmetic genus two with ample canonical bundle. Then the following hold:
  1. (i)
    The bicanonical map is a morphism of degree two onto a conic $Q$ in ${\mathbb{P}}^{2}$.
  2. (ii)
    The conic $Q$ is reducible if and only if $C$ is of type B.
  3. (iii)
    The rational involution $\iota$ induced on $C$ by the bicanonical map is regular.
Proof.
$(i)$ and $(ii)$ follow immediately by Proposition 3.3.
$(iii)$ By Proposition 3.3 the curve $C$ is embedded in a weighted projective space $\mathbb{P}$ in such a way that the regular involution of $\mathbb{P}$ that sends $u$ to $-u$ and leaves the remaining variables invariant induces $\iota$ on $C$. $\square$
Remark 3.7
Let $X$ be a variety (so in particular a Cohen–Macaulay scheme, cf. section 2.A) with an involution $\iota$ that acts non-trivially on every irreducible component of $X$ and let $\pi\colon X\to Y:=X/\iota$ be the quotient map. Then $\pi_{*}{\mathcal{O}}_{X}$ decomposes under the action of $\iota$ as a direct sum ${\mathcal{O}}_{Y}\oplus{\mathcal{F}}$, where ${\mathcal{F}}$ is a torsion-free sheaf of rank one, and $\iota$ acts on ${\mathcal{O}}_{Y}$ as the identity and on ${\mathcal{F}}$ as multiplication by $-1$. By relative duality we have $\pi_{*}\omega_{X}=\curly{H}om(\pi_{*}{\mathcal{O}}_{X},\omega_{Y})$ and the decomposition of $\pi_{*}{\mathcal{O}}_{X}$ induces a decomposition $\pi_{*}\omega_{X}=\omega_{Y}\oplus\curly{H}om({\mathcal{F}},\omega_{Y})$, where the first summand is $\iota$-invariant and the second one is anti-invariant. We call the corresponding linearisation of $\omega_{Y}$ “natural”. Note that when both $X$ and $Y$ are smooth the natural linearisation corresponds to the natural action on differential forms, for which invariant $n$-forms on $X$ correspond to pull-backs of $n$-forms on $Y$; note also that if the map $\pi$ is flat, namely ${\mathcal{F}}$ is a line bundle, then $\omega_{X}=\varphi^{*}(\omega_{Y}\otimes{\mathcal{F}}^{-1})$ and the opposite of the natural linearisation is the one induced by the fact that $\omega_{X}$ is a pull-back from $Y$.
Consider now a reduced Gorenstein curve $C$ of genus two with $\omega_{C}$ ample; the bicanonical map $\varphi\colon C\to Q$ is the quotient map for the bicanonical involution $\iota$ defined in Corollary 3.6. Actually in Corollary 3.6 $\iota$ is defined by its action on the canonical ring $R(C,\omega_{C})$, which in addition determines a linearisation of $\omega_{C}$ for which the elements of $H^{0}(\omega_{C})$ are all invariant. Since $H^{0}(\omega_{Q})=0$, this is the opposite of the natural linearisation and is the linearisation that we will always consider in our computations
Corollary 3.8 —
With respect to the linearisation fixed in Remark 3.7, the invariant subring $R(C,\omega_{C})^{+}\subset R(C,\omega_{C})$ has Hilbert series $(1-t^{2})^{-2}$.
Proof.
The invariant subring $R(C,\omega_{C})^{+}\subset R(C,\omega_{C})$ is the ring generated by $y_{1},y_{2}$ for type A and by $y_{1},y_{2},v$ for type B, so its Hilbert series is $\frac{1}{(1-t^{2})^{2}}=\frac{1-t^{4}}{(1-t^{2})^{2}(1-t^{4})}$. $\square$
We close this section by describing the geometry of the curves described in Proposition 3.3. The possibilities are illustrated in Table 1.
Corollary 3.9 —
Let $C$ be a reduced connected Gorenstein curve of arithmetic genus two with ample canonical bundle. Then
  1. (i)
    If $C$ is of type A then:
    • (a)
      the singularities of $C$ are of type $A_{n_{1}},\dots A_{n_{k}}$, with $\sum_{i}(n_{i}+1)\leq 6$;
    • (b)
      $C$ is reducible iff $\sum_{i}(n_{i}+1)=6$ and all the $n_{i}$ are odd. The components of $C$ are smooth rational curves exchanged by $\iota$.
  2. (ii)
    If $C$ is of type $B$, then it is the union of two irreducible curves $C_{i}$ with $p_{a}(C_{i})=1$ for $i=1,2$, that meet transversely at a single point $p$ that is smooth for both. The involution $\iota$ restricts on each component to an involution with smooth rational quotient curve.
Proof.
$(i)$ Up to a linear change in the coordinates $y_{1},y_{2}$ we may assume that an open set $U$ of $C$ containing all the singularities is $U=\{u^{2}-\Pi_{1}^{r}(y-\alpha_{i})^{m_{i}}=0\}\subset\mathbb{A}^{2}_{y,u}$. where the $\alpha_{i}\in\mathbb{C}$ are distinct, $\sum m_{i}=6$, and moreover $m_{i}$ is odd if and only if $i\leq k$, namely we group all the odd $m_{i}$ at the beginning. The singularities are the points $(\alpha_{i},0)$ with $m_{i}>1$ and they are of type $A_{n_{i}}$, $n_{i}:=m_{i}-1$. The normalization of $U$ is $u^{2}-\Pi_{1}^{k}(y-\alpha_{i})$, so it is reducible if and only if all the $m_{i}$ are even.
$(ii)$ We have already noticed in the proof of Proposition 3.3 that $C_{1}$ and $C_{2}$ have arithmetic genus 1 and meet in a point $p$ that is a node for $C$. The rest is an immediate consequence of Proposition 3.3 and Corollary 3.6. $\square$
Table 1. The possible curves of type A and B, with components marked with the geometric genus
Type A Type B
4. Generalised Gorenstein spin curves
A spin structure (or theta-characteristic) on a smooth curve is a line bundle ${\mathcal{L}}$ such that there is an isomorphism $\mu\colon{\mathcal{L}}\otimes{\mathcal{L}}\to\omega_{C}$. Here we propose a generalisation of this notion to Gorenstein curves. We remark that in the case of (semi-)stable curves there are several analogous notions in the literature (see e.g., [Cor89, Jar98, Lud10]).
Definition 4.1 —
Let $C$ be a projective Gorenstein curve. A generalised Gorenstein spin (ggs) structure on $C$ is a pair $({\mathcal{L}},\mu)$ where ${\mathcal{L}}$ is a torsion-free sheaf of rank $1$ on $C$ with $\chi({\mathcal{L}})=0$, and
$\mu\colon{\mathcal{L}}\operatorname{\otimes}_{{\mathcal{O}}_{C}}{\mathcal{L}}\to{\mathcal{L}}^{[{2}]}\to\omega_{C}$
is a map which is an isomorphism in codimension zero, that is, at the generic point of each component of $C$. We say the ggs structure is regular if ${\mathcal{L}}$ is a line bundle.
Isomorphisms of ggs structures are defined in the obvious way, namely, ggs structures $(C_{1},{\mathcal{L}}_{1},\mu_{1})$ and $(C_{1},{\mathcal{L}}_{1},\mu_{1})$ are isomorphic if there are an isomorphism $\varphi\colon C_{1}\to C_{2}$ and an isomorphism ${\mathcal{L}}_{1}\to\varphi^{*}{\mathcal{L}}_{2}$ compatible with $\mu_{1}$ and $\varphi^{*}\mu_{2}$. In particular, when $\varphi$ is the identity we may rescale $\mu$ by a non-zero scalar.
Note that the second condition implies that ${\mathcal{L}}$ is generically locally free, since $\omega_{C}$ is locally free. This is automatic on a reduced curve.
Note also that for reasons of degree the multiplication map is an isomorphism if the ggs structure is regular. If $C$ is smooth, all ggs structures are regular because a torsion-free sheaf is locally free at a smooth point of any curve.
In some applications ggs structures will appear naturally also on non-reduced curves but uniform results seem to be much harder to come by, so we do not address this case here.
Let $C$ be a reduced curve and $f\colon\widetilde{C}\to C$ a finite birational morphism. Then the normalisation of $C$ factors over $f$, which is therefore often called a partial normalisation. Note that a partial normalisation might not preserve the Gorenstein property in general: three general lines through the origin in ${\mathbb{C}}^{3}$ are not Gorenstein, but map birationally to three general lines through the origin in the plane, a Gorenstein singularity.
Example 4.2 —
Let us describe all partial normalisations locally at a point of type $A_{m}$ where we have two cases according to the parity of $m$, compare [Yos90, 9.9, 5.11].
$A_{2n}$:
The completion of the local ring is isomorphic to the ring ${\mathbb{C}}\llbracket x,y\rrbracket/(y^{2}-x^{2n+1})\cong{\mathbb{C}}\llbracket t^{2},t^{2n+1}\rrbracket$ and the partial normalisations are given by the inclusions
${\mathbb{C}}\llbracket t\rrbracket\supset{\mathbb{C}}\llbracket t^{2},t^{3}\rrbracket\supset{\mathbb{C}}\llbracket t^{2},t^{5}\rrbracket\supset{\mathbb{C}}\llbracket t^{2},t^{7}\rrbracket\supset\dots$
$A_{2n-1}$:
The completion of the local ring is isomorphic to ${\mathbb{C}}\llbracket x,y\rrbracket/(y^{2}-x^{2n})$ and we get a series of inclusions
${\mathbb{C}}\llbracket x,y\rrbracket/(y-x)\times{\mathbb{C}}\llbracket x,y\rrbracket/(y+x)\hookleftarrow{\mathbb{C}}\llbracket x,y\rrbracket/(y^{2}-x^{2})\supset{\mathbb{C}}\llbracket x,xy\rrbracket\supset{\mathbb{C}}\llbracket x,x^{2}y\rrbracket\supset\dots$
where we consider ${\mathbb{C}}\llbracket x,x^{k}y\rrbracket$ as a subring of ${\mathbb{C}}\llbracket x,y\rrbracket/(y^{2}-x^{2})$, so that with $y_{k}=x^{k}y$ we get the desired relation $y_{k}^{2}-x^{k+2}=0$.
Note that in both series each successive subring has codimension one in the previous ring, when considered as a ${\mathbb{C}}$-vector space.
From this description we see that any partial normalisation of an $A_{n}$ singularity is an $A_{m}$ singularity with $m<n$ of the same parity, or smooth. In particular, it is also Gorenstein.
The following is true if every partial normalisation of the Gorenstein curve $C$ is again Gorenstein, but we state it only for the case that we need.
Proposition 4.3 —
Let $({\mathcal{L}},\mu)$ be a ggs structure on a reduced Gorenstein curve with only singularities of type $A_{m}$. Then there exists a (uniquely determined) partial normalisation $f\colon\widetilde{C}\to C$ and a regular ggs structure (~$(\widetilde{\mathcal{L}},\widetilde{\mu})$ on $\widetilde{C}$ such that
${\mathcal{L}}=f_{*}\widetilde{\mathcal{L}}\text{ and }{\mathcal{L}}^{[{2}]}=f_{*}\widetilde{\mathcal{L}}^{\operatorname{\otimes}2}.$
Proof.
Since at a smooth point of a curve every torsion-free sheaf is free, the question is local at the singular points of $C$. Then the existence of the partial normalisation and the line bundle $\widetilde{\mathcal{L}}$ follows from Lemma 2.3 its unicity from the description in Example 4.2.
For the last statement, fix a singular point $P\in C$ and denote by $Z$ its preimage in $\widetilde{C}$. Note that it is possible to find an affine open set of $\widetilde{C}$ containing $Z$ on which $\widetilde{{\mathcal{L}}}$ is trivial. In fact, twisting $\widetilde{{\mathcal{L}}}$ with a sufficiently ample effective divisor $H$ disjoint from $Z$ we may assume that $|{\mathcal{L}}(H)|$ is base point free. So there is a section $\sigma\in H^{0}(\widetilde{\mathcal{L}}(H))$ that does not vanish at any point of $Z$, and therefore trivializes $\widetilde{\mathcal{L}}$ on an affine open set containing $Z$ and contained in the complement of $H$. Therefore we may replace $C$ by affine open neighbourhood $U=\operatorname{\rm Spec}A$ of $P$, $\widetilde{C}$ by the preimage $\widetilde{U}=\operatorname{\rm Spec}\widetilde{A}$ of $U$ in $\widetilde{C}$ and $\widetilde{\mathcal{L}}$ by the trivial bundle. The map $f_{*}\widetilde{\mathcal{L}}\otimes f_{*}\widetilde{\mathcal{L}}\to f_{*}\widetilde{\mathcal{L}}^{\otimes 2}$ corresponds to the map of $A$-modules $\widetilde{A}\otimes_{A}\widetilde{A}\to\widetilde{A}$ induced by the multiplication of $\widetilde{A}$, and thus it is surjective. Since both sheaves have rank one and $f_{*}\widetilde{\mathcal{L}}$ is torsion-free, the induced map ${\mathcal{L}}^{[{2}]}\to f_{*}\widetilde{\mathcal{L}}^{\operatorname{\otimes}2}$ is an isomorphism.
It remains to show that (~$(\widetilde{\mathcal{L}},\widetilde{\mu})$ is a regular ggs structure on $\widetilde{C}$. We have seen in Example 4.2 that $\widetilde{C}$ is again Gorenstein, so we can apply relative duality for a finite morphism of Gorenstein schemes, e.g., [Sta23, Section 0AU3], which gives
$\operatorname{Hom}({\mathcal{L}}^{[2]},\omega_{C})=\operatorname{Hom}(f_{*}\widetilde{\mathcal{L}}^{\operatorname{\otimes}2},\omega_{C})\cong\operatorname{Hom}(\widetilde{\mathcal{L}}^{\operatorname{\otimes}2},\omega_{\widetilde{C}}),$
and clearly $0=\chi({\mathcal{L}})=\chi(\widetilde{\mathcal{L}})$. Restriction to the subset where $f$ is an isomorphism shows that $\varphi\colon{\mathcal{L}}^{[{2}]}\to\omega_{C}$ is generically an isomorphism if and only if the corresponding $\widetilde{\varphi}\colon\widetilde{\mathcal{L}}^{\operatorname{\otimes}2}\to\omega_{\widetilde{C}}$ is, completing the correspondence between ggs structures. $\square$
With Proposition 4.3 in place, the enumeration of regular and irregular ggs structures on any given curve with type $A_{n}$ singularities is relatively simple, but the number of cases to consider grows quickly, if there are several possible partial normalisations to consider.
Example 4.4 —
On a smooth rational curve, the only ggs structure is ${\mathcal{L}}\cong{\mathcal{O}}_{{\mathbb{P}}^{1}}(-1)$.
Proposition 4.5 —
Let $C=\bigcup_{i}C_{i}$ be a reduced and connected Gorenstein curve, where $C_{i}$ are the irreducible components.
  1. (i)
    The curve $C$ admits a regular ggs structure $({\mathcal{L}},\mu)$ if and only if $\deg\omega_{C}|_{C_{i}}$ is even on every component $C_{i}$.
  2. (ii)
    If $C$ admits a regular ggs structure, then it admits exactly $2^{b_{1}}$ different regular ggs structures, where $b_{1}$ is the first Betti number of $C$.
In particular, every irreducible Gorenstein curve admits a regular ggs structure while a curve with a separating node does not admit a regular ggs structure .
Proof.
Let $({\mathcal{L}},\mu)$ be a regular ggs structure. Then for any $2$-torsion line bundle $\eta\in\operatorname{Pic}^{0}(C)$ the bundle ${\mathcal{L}}\operatorname{\otimes}\eta$ is a again a ggs structure and the set of regular ggs structures is a torsor under the group of $2$-torsion line bundles $\operatorname{Pic}^{0}(C)[2]$. By the exponential sequence, $\operatorname{Pic}^{0}(C)$ is a connected algebraic group that fits in an exact sequence
(4.6) $0\to H^{1}(C,{\mathbb{Z}})\to H^{1}(C,{\mathcal{O}}_{C})\to\operatorname{Pic}^{0}(C)\to 0$
and therefore $2$-torsion elements in this group are given by the image of $\frac{1}{2}H^{1}(C,{\mathbb{Z}})$ in $\operatorname{Pic}^{0}(C)$. This proves $(ii)$.
If $\omega_{C}|_{C_{i}}$ has odd degree on any component, then there cannot exist a regular ggs structure $({\mathcal{L}},\mu)$ on $C$, because $2\deg{\mathcal{L}}|_{C_{i}}=\deg\omega_{C}|_{C_{i}}$.
On the other hand, assume that $\deg\omega_{C}|_{C_{i}}=2d_{i}$ for all $i$ and pick a smooth point $p_{i}$ in $C_{i}$. Then $\eta:=\omega_{C}\left(-\sum_{i}2d_{i}p_{i}\right)$ has degree zero on each component and thus $\eta\in\operatorname{Pic}^{0}(C)$. Since the group $\operatorname{Pic}^{0}(C)$ is 2-divisible, for example by the description in (4.6), we can pick an element $\eta^{\prime}$ such that $\eta^{\prime\operatorname{\otimes}2}=\eta$. Then $\eta^{\prime}\left(\sum_{i}d_{i}p_{i}\right)$ is a regular ggs structure on $C$. $\square$
We now collect the necessary information about ggs structures on the curves we are interested in.
Proposition 4.7 —
Let $C$ be a Gorenstein curve of arithmetic genus two with $\omega_{C}$ ample, bicanonical involution $\iota$ and ggs structure $({\mathcal{L}},\mu)$. Then
  1. (i)
    ${\mathcal{L}}$ is invariant under the involution, that is, $\iota^{*}{\mathcal{L}}\cong{\mathcal{L}}$;
  2. (ii)
    $h^{0}(C,{\mathcal{L}})\leq 2$ and equality holds if and only if $C=C_{1}\cup C_{2}$ is of Type $B$ and ${\mathcal{L}}={\mathcal{O}}_{C_{1}}\oplus{\mathcal{O}}_{C_{2}}$ is the pushforward of the unique ggs structure with a non-zero section on each component.
  3. (iii)
    $\mu\colon{\mathcal{L}}\operatorname{\otimes}{\mathcal{L}}\to\omega_{C}$ is an isomorphism on the locus where ${\mathcal{L}}$ is invertible.
Proof.
Let us write ${\mathcal{L}}=f_{*}\widetilde{\mathcal{L}}$ for a regular ggs structure (~$(\widetilde{\mathcal{L}},\widetilde{\mu})$ on a partial normalisation $f\colon\widetilde{C}\to C$. Recall that by Corollary 3.9 the curve $\widetilde{C}$ is also Gorenstein. The composition $f\circ\varphi:\widetilde{C}\to Q$ is a degree 2 finite morphism, so it induces an involution $\widetilde{\iota}$ on $\widetilde{C}$ that lifts $\iota$.
We first observe that $\iota$ acts trivially on $\operatorname{Pic}^{0}(\widetilde{C})[2]$. Indeed, considering the induced action on $H^{1}({\mathcal{O}}_{\widetilde{C}})\cong H^{0}(\widetilde{C},\omega_{\widetilde{C}})$ and taking into account that $\widetilde{C}/\widetilde{\iota}$ has arithmetic genus 00 we see $\widetilde{\iota}$ acts on $H^{1}(\widetilde{C},{\mathcal{O}}_{\widetilde{C}})$ as multiplication by $-1$. Then the claim follows from the exponential sequence. It thus suffices to exhibit one regular $\widetilde{\iota}$-invariant ggs structure.
By the criterion in Proposition 4.5 $(i)$ the curve $\widetilde{C}$ admits a regular ggs structure if and only if it is irreducible or $\deg\omega_{C}|_{C_{i}}$ is even on both components.
So if $C$ is of type B then $\widetilde{C}$ is disconnected and its connected components are irreducible of genus zero or one (cf. Table 1), so we can take the trivial bundle on components of genus one and the unique ggs structure on components of genus zero and the resulting ggs structure is obviously invariant.
Consider now $C$ of type A. By Corollary 3.9 if $\widetilde{C}$ is disconnected then it is the union of two smooth rational curves and we can take the unique ggs structure on each of them. If $\widetilde{C}$ is connected but reducible, then it is the union of two smooth rational curves, meeting either in two nodes or in an $A_{3}$ point, so it has genus 1 and we can take the trivial ggs structure ${\mathcal{O}}_{\widetilde{C}}$ on $\widetilde{C}$. If $\widetilde{C}\neq C$ is irreducible, then it has arithmetic genus at most one and admits either a trivial or a unique ggs structure, as above.
So we are left with the case when $C=\widetilde{C}$ is irreducible. If $\varphi\colon C\to C/\iota={\mathbb{P}}^{1}$ admits a smooth branch point $p$, then the associated line bundle is a regular ggs structure, since $\omega_{C}=\varphi^{*}{\mathcal{O}}_{{\mathbb{P}}^{1}}(1)={\mathcal{O}}_{C}(2p)$. The only remaining case is when $\widetilde{C}$ has two points of type $A_{2}$. In this case, $b_{1}(C)=0$ so by Proposition 4.5 there exists a unique ggs structure on $C$, which then has to be $\iota$-invariant.
We move on to the second item. First observe that if $C$ is a connected Gorenstein curve of genus $g\leq 2$, then for any regular ggs structure ${\mathcal{L}}$ on $C$ we have $h^{0}(C,{\mathcal{L}})\leq 1$, because $\deg{\mathcal{L}}=g-1$.
If $g=2$ and $C$ is reducible, then by Proposition 4.5 and the geometry of the curves (Proposition 3.3) any ggs structure is irregular, so we consider the corresponding regular ggs structure (~$(\widetilde{\mathcal{L}},\widetilde{\mu})$ on a partial normalisation $\widetilde{C}$. If $\widetilde{C}$ is connected, then $h^{0}(\widetilde{\mathcal{L}})\leq 1$ by the above. If $\widetilde{C}$ is disconnected, then both components have genus at most one, so $h^{0}(C,{\mathcal{L}})\leq 2$ with equality if and only if $C=C_{1}\cup C_{2}$ is of type $B$ and ${\mathcal{L}}={\mathcal{O}}_{C_{1}}\oplus{\mathcal{O}}_{C_{2}}$ is the pushforward of the unique ggs structure with a section on the components, because the unique regular ggs structure on ${\mathbb{P}}^{1}$ does not have a section.
$(iii)$ Follows from the fact that the points where ${\mathcal{L}}$ is invertible are those where $\widetilde{C}\to C$ is an isomorphism and from the fact that $\mu$ is an isomorphism for regular ggs structures. $\square$
We close this section with a technical result needed for the computations of Section 5.
Lemma 4.8 —
Let $C$ be a reduced connected Gorenstein curve of arithmetic genus two with $\omega_{C}$ ample and ggs structure $({\mathcal{L}},\mu)$. Assume that $C$ is reducible and let $B\subset C$ be an irreducible component. Then $\deg({\mathcal{L}}{\raisebox{-0.1pt}{$|_{B}$}}^{[1]})\geq-1$ and if equality holds then $p_{a}(B)=0$.
Proof.
By Proposition 4.3 we have ${\mathcal{L}}=f_{*}\widetilde{\mathcal{L}}$ for a regular ggs structure $\widetilde{\mathcal{L}}$ on a partial normalisation $f\colon\widetilde{C}\to C$. Denote by $\widetilde{B}$ the component of $\widetilde{C}$ that maps to $B$ and denote by ${\mathcal{I}}\subset{\mathcal{O}}_{\widetilde{C}}$ the ideal of $\widetilde{B}$. Since $f$ is finite, we have an exact sequence
(4.9) $0\to f_{*}({\mathcal{I}}\widetilde{\mathcal{L}})\to f_{*}\widetilde{\mathcal{L}}={\mathcal{L}}\to f_{*}(\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}})\to 0$
The sheaf $f_{*}(\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}})$ is torsion-free and supported on $B$, so the last map in the above equation gives a surjective map ${\mathcal{L}}{\raisebox{-0.1pt}{$|_{B}$}}^{[1]}\to f_{*}(\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}})$. Since both sheaves are torsion-free of rank one, the map is an isomorphism.
So we have $\chi({\mathcal{L}}{\raisebox{-0.1pt}{$|_{B}$}}^{[1]})=\chi({\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}})=\deg\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}}+1-p_{a}(\widetilde{B}})$ and $2\deg(\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}})=\deg K_{\widetilde{C}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}}$. Since there is an inclusion $\omega_{\widetilde{B}}\hookrightarrow\omega_{\widetilde{C}}$ (cf. Lemma 2.4), we get $\deg\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}}\geq p_{a}(\widetilde{B})-1\geq-1$, and if $\deg\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}}=-1$ holds then $\widetilde{B}$ is smooth rational.
Assume $\deg\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}}=-1$, hence $0=\chi(\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}})=\chi({\mathcal{L}}{\raisebox{-0.1pt}{$|_{B}$}}^{[1]})$: if $B$ is not smooth rational then $p_{a}(B)=1$ and $\deg{\mathcal{L}}{\raisebox{-0.1pt}{$|_{B}$}}^{[1]}=0$. $\square$
Corollary 4.10 —
Let $C$ be a reduced connected Gorenstein curve of arithmetic genus two with $\omega_{C}$ ample and ggs structure $({\mathcal{L}},\mu)$. Then we have $H^{1}(C,{\mathcal{L}}\otimes\omega_{C}^{\otimes m})=0$ for every $m\geq 1$.
Proof.
We have $\deg({\mathcal{L}}\otimes\omega_{C}^{\otimes m})=2m+1\geq 3=2p_{a}(C)-1$, since $\omega_{C}$ is a line bundle of degree two. Therefore if $C$ is irreducible we can conclude by Lemma 2.5.
Assume now that $C$ is reducible and let $B\subset C$ be an irreducible component, so that $\omega_{C}{\raisebox{-0.1pt}{$|_{B}$}}$ is a line bundle of degree one. Then by the above Lemma 4.8 we have $\deg(({\mathcal{L}}\otimes\omega_{C}^{\otimes m}){\raisebox{-0.1pt}{$|_{B}$}}^{[1]})\geq m-1$ if $p_{a}(B)=0$ and $\geq m$ if $p_{a}(B)=1$. Summing up, we have $\deg({\mathcal{L}}\otimes\omega_{C}^{\otimes m}){\raisebox{-0.1pt}{$|_{B}$}}^{[1]}\geq 2p_{a}(B)-1$ for every $B\subseteq C$, hence Lemma 2.5 gives $H^{1}({\mathcal{L}}\otimes\omega_{C})=0$. $\square$
5. Half-canonical rings
5.A. Definition and classification of half-canonical rings
Definition 5.1 —
Let $C$ be a projective and reduced Gorenstein curve and let $({\mathcal{L}},\mu)$ be a ggs structure on $C$. Its half-canonical ring is
$R(C,\{{\mathcal{L}},\omega_{C}\})=R(C,{\mathcal{L}})=\bigoplus_{n}R_{n},$
where $R_{2n}=H^{0}(C,nK_{C})$ and $R_{2n+1}=H^{0}(C,{\mathcal{L}}(nK_{C}))$ and the multiplication of two elements of odd degree is defined via $\mu$.
Now let $({\mathcal{L}},\mu)$ be a ggs structure on a curve $C$ with ample canonical bundle and arithmetic genus two. We classify the possible half-canonical rings $R(C,\{{\mathcal{L}},\omega_{C}\})$ according to the type of the curve defined in Proposition 3.3.
Theorem 5.2 —
Let $({\mathcal{L}},\mu)$ be a ggs structure on a curve $C$ of genus two with ample canonical bundle. Then $h^{0}({\mathcal{L}})=0,1,2$ and the half-canonical ring $R(C,\{{\mathcal{L}},\omega_{C}\})$ admits generators and relations as given in Table 2.
Conversely, under the assumptions on $f_{6}$, $g_{6}$, $g_{8}$, $f_{12}$ shown in Table 2, the ideals of Table 2 define reduced Gorenstein curves $C$ of genus two with $\omega_{C}={\mathcal{O}}_{C}(2)$ ample and ${\mathcal{L}}={\mathcal{O}}_{C}(1)^{[{1}]}$ a ggs structure on $C$.
Table 2. Half-canonical rings in terms of generators and relations
$h^{0}({\mathcal{L}})$ Type variables & degrees relations
0 $A(0)$ $\begin{pmatrix}y_{1}&y_{2}&z_{1}&z_{2}\\ 2&2&3&3\end{pmatrix}$ $z_{1}^{2}-f_{6}(y_{1},y_{2}),\,z_{2}^{2}-g_{6}(y_{1},y_{2})$
where $f_{6}\neq 0$, $g_{6}\neq 0$
$B(0)$ $\begin{pmatrix}y_{1}&y_{2}&z_{1}&z_{2}&v&t_{1}&t_{2}&u\\ 2&2&3&3&4&5&5&6\end{pmatrix}$ $\operatorname{rk}\begin{pmatrix}y_{1}&0&z_{1}&t_{1}\\ 0&y_{2}&z_{2}&t_{2}\\ z_{1}&z_{2}&v&u\\ t_{1}&t_{2}&u&g_{8}\end{pmatrix}\leq 1$ where $g_{8}=g_{8}(y_{1},y_{2},v)$ and $v^{2}$ appears in $g_{8}$ with non-zero coefficient.
1 $A(1)$ $\begin{pmatrix}x&y&z\\ 1&2&5\end{pmatrix}$ $z^{2}-f_{10}(x,y)$
where $f_{10}\neq 0$
$B(1)$ $\begin{pmatrix}x&y&w&v&z&u\\ 1&2&3&4&5&6\end{pmatrix}$ $\operatorname{rk}\begin{pmatrix}0&y&w&z\\ x&w&v&u\end{pmatrix}\leq 1,$ $\begin{array}[]{rcl}z^{2}&=&yg_{8}(y,v)\\ zu&=&wg_{8}(y,v)\\ u^{2}&=&vg_{8}(y,v)+x^{4}h_{8}(x,v)\end{array}$ where $v^{2}$ appears in $g_{8}$ with non-zero coefficient.
2 $B(2)$ $\begin{pmatrix}x_{1}&x_{2}&v&u\\ 1&1&4&6\end{pmatrix}$ $x_{1}x_{2},\ \ u^{2}-f_{12}(x_{1},x_{2},v)$ where $v^{3}$ appears in $f_{12}$ with non-zero coefficient.
Remark 5.3
We have the following geometric interpretations of the type A rings:
  1. (i)
    If $h^{0}({\mathcal{L}})=1$ then $C$ is a double cover of ${\mathbb{P}}^{1}$ with coordinates $x^{2},y$ with ramification divisor $R$ consisting of the five points determined by $f_{10}$ together with the point $p$ defined by $x^{2}=0$. If $C$ is smooth at $p$, then ${\mathcal{L}}={\mathcal{O}}_{C}(p)$.
  2. (ii)
    If $h^{0}({\mathcal{L}})=0$ then $C$ is a double cover of ${\mathbb{P}}^{1}$ with coordinates $y_{1},y_{2}$ with ramification divisor $R$ consisting of the three points $p_{1},p_{2},p_{3}$ determined by $f_{6}$ and the three points $q_{1},q_{2},q_{3}$ determined by $g_{6}$. If $C$ is smooth, then ${\mathcal{L}}={\mathcal{O}}(p_{1}+p_{2}-p_{3})={\mathcal{O}}(q_{1}+q_{2}-q_{3})$.
In both cases, singularities of $C$ correspond to multiple points of the branch locus.
Remark 5.4
The type B curves $C=C_{1}\cup C_{2}$ are double covers of the reducible conic $Q_{1}\cup Q_{2}$ branched on the point $Q_{1}\cap Q_{2}$ and on three more points on each component, cf. Corollary 3.6. The geometry of the half-canonical rings is as follows:
  1. (i)
    If $h^{0}({\mathcal{L}})=0$ then one component is a double cover of ${\mathbb{P}}^{1}$ with coordinates $y_{1}^{2},v$ branched over the points $v=0$ and $y_{1}=0$ and the two points determined by $g_{8}|_{y_{2}=0}$. The other component is a double cover of ${\mathbb{P}}^{1}$ with coordinates $y_{2}^{2},v$ branched on the points $v=0$ and $y_{2}=0$and the two points determined by $g_{8}|_{y_{1}=0}$.
  2. (ii)
    If $h^{0}({\mathcal{L}})=1$ then one component is a double cover of ${\mathbb{P}}^{1}$ with coordinates $x^{4},v$ branched in $[0,1]$ and in the three points determined by $v^{3}+x^{4}h_{8}$. On the other component, the double cover of ${\mathbb{P}}^{1}$ can be viewed in coordinates $y^{2},v$, branched in the point $v=0$ and the two points determined by $g_{8}$.
  3. (iii)
    If $h^{0}({\mathcal{L}})=2$ then one component is a double cover of ${\mathbb{P}}^{1}$ with coordinates $x_{1}^{4},v$ branched in $[0,1]$ and in the three points determined by $h_{12}|_{x_{2}=0}$. The other component is a double cover of ${\mathbb{P}}^{1}$ with coordinates $x_{2}^{4},v$ branched in in $[0,1]$ and the three points determined by $h_{12}|_{x_{1}=0}$.
We prove Theorem 5.2 case-by-case, starting from the description of the canonical ring
$R(C,\omega_{C})\subset R=R(C,\{{\mathcal{L}},\omega_{C}\})$
obtained in Proposition 3.3 and, for type B, exploiting the involution $\iota$ on $C$ induced by the bicanonical map.
5.B. First observations
We begin by showing that the Hilbert series of $R$ is determined by the value of $h^{0}({\mathcal{L}})$.
Proposition 5.5 —
Let $({\mathcal{L}},\mu)$ be a ggs structure on a curve of genus two with ample canonical bundle. Then the Hilbert series of the half-canonical ring $R$ is
$h^{0}(C,{\mathcal{L}})$ Hilbert series of $R$
0 $\frac{(1-t^{6})(1-t^{6})}{(1-t^{2})(1-t^{2})(1-t^{3})(1-t^{3})}$
1 $\frac{1-t^{10}}{(1-t)(1-t^{2})(1-t^{5})}$
2 $\frac{(1-t^{2})(1-t^{12})}{(1-t)(1-t)(1-t^{4})(1-t^{6})}$
Proof.
The dimension of $R_{2}=H^{0}(K_{C})$ is two and for $m>1$ we know $\dim R_{2m}=h^{0}(C,mK_{C})=2m-1$ from Proposition 3.3, so we only need to determine the dimension of $R_{2m+1}=H^{0}({\mathcal{L}}(mK_{C}))$.
The dimension of $R_{1}=H^{0}({\mathcal{L}})$ is fixed by our assumption. Concerning the other odd degrees, by Corollary 4.10 we have $H^{1}(C,{\mathcal{L}}(mK_{C}))=0$, whence by Riemann–Roch we obtain $h^{0}(C,{\mathcal{L}}(mK_{C}))=2m$ for every $m\geq 1$.
Putting this all together as in the proof of Proposition 3.3 gives the stated Hilbert series for the ring $R$. $\square$
The next result is used to bound our computations of half-canonical rings.
Lemma 5.6 —
  1. (i)
    If $({\mathcal{L}},\mu)$ is a ggs structure on a curve of Type A then $R(C,\{{\mathcal{L}},\omega_{C}\})$ is generated in degrees $\leq 6$.
  2. (ii)
    If $({\mathcal{L}},\mu)$ is a ggs structure on a curve of Type B then $R(C,\{{\mathcal{L}},\omega_{C}\})$ is generated in degrees $\leq 9$.
Proof.
$(i)$ By Proposition 3.3, if $C$ is of Type A, then $|K_{C}|$ is base point free. Hence by Proposition 2.6 and Corollary 4.10 the following multiplication maps
$H^{0}(K_{C})\operatorname{\otimes}H^{0}(mK_{C})\to H^{0}((m+1)K_{C})=R_{2m+2},$
$\ H^{0}(K_{C})\operatorname{\otimes}H^{0}\left({\mathcal{L}}\left((m-1)\right)K_{C}\right)\to H^{0}({\mathcal{L}}(mK_{C}))=R_{2m+1}$
are surjective for $m\geq 3$. Hence $R$ is generated in degrees $\leq 6$.
$(ii)$ Since the linear system $|2K_{C}|$ is base point free and since for $m\geq 3$ Corollary 4.10 gives $H^{1}(C,{\mathcal{L}}((m-2)K_{C}))=0$, Proposition 2.6 applies to ${\mathcal{L}}(mK_{C})$ and $2K_{C}$ and so the multiplication map
$H^{0}(2K_{C})\operatorname{\otimes}H^{0}({\mathcal{L}}(mK_{C}))\to H^{0}({\mathcal{L}}((m+2)K_{C}))=R_{2m+5}$
is surjective for all $m\geq 3$. It follows that $R$ is generated in degrees $\leq 10$. Since the canonical ring is generated in degrees $\leq 6$ (Proposition 3.3), we also know that there are no new generators in degree $10$square
5.C. Type A
Let $C$ be a curve of type A, i.e., by Corollary 3.9, $C$ is irreducible or $C=C_{1}\cup C_{2}$ with $p_{a}(C_{i})=0$. We say that a ggs structure $({\mathcal{L}},\mu)$ is of type $A(i)$ if $C$ is of type $A$ and $h^{0}(C,{\mathcal{L}})=i$. By Proposition 4.7 we have $i=0$ or $i=1$.
We are now going to prove cases $A(1)$ and $A(0)$ of Theorem 5.2.
5.C.1. Type $A(1)$
Here we prove that if $C$ is of type $A(1)$ then $R(C,\{{\mathcal{L}},\omega_{C}\})$ is isomorphic to ${\mathbb{C}}[x,y,z]/(z^{2}-f_{10})$, with $\deg(x,y,z)=(1,2,5)$.
Proof.
Choose a non-zero section $x\in R_{1}=H^{0}({\mathcal{L}})$. Since $C$ is reduced we have $x^{2}\neq 0$ and thus we can choose the coordinates in the canonical ring such that $x^{2}=y_{1}$ and let $y:=y_{2}$. We start by noting the following facts:
  • (a)
    Multiplication by $x$ gives an injective map $R(C,\{{\mathcal{L}},\omega_{C}\})\to R(C,\{{\mathcal{L}},\omega_{C}\})$. Indeed, we have seen in the proof of Proposition 3.3 that a curve of type A either is irreducible, or it is the union of two smooth rational curves $C_{1}$ and $C_{2}$ and the restriction map $H^{0}(\omega_{C})\to H^{0}(\omega_{C}{\raisebox{-0.1pt}{$|_{C_{i}}$}})=H^{0}(\mathcal{O}_{C_{i}}(1))$ is an isomorphism for $i=1,2$. So $x^{2}$, and a fortiori $x$, does not vanish at the generic point of any component of $C$.
  • (b)
    The map $j\colon{\mathbb{C}}[x,y]\to R(C,\{{\mathcal{L}},\omega_{C}\})$ is injective. Let $v$ be a nonzero element of $\ker j$: since $j$ is a map of graded rings, we may assume that $v$ is homogeneous. Then $\deg v$ is odd, since the map ${\mathbb{C}}[x^{2},y]\to R(C,\omega_{C})$ is an inclusion by Proposition 3.3 and we can write $v=xv_{1}$ with $v_{1}$ homogeneous of even degree. By (a) $v_{1}$ is in the kernel of ${\mathbb{C}}[x^{2},y]\to R(C,\omega_{C})$, a contradiction.
In view of fact (b) above, comparing dimensions we see that ${\mathbb{C}}[x,y]_{\leq 4}=R_{\leq 4}$ and that
$R_{5}=\langle x^{5},x^{3}y,xy^{2},z\rangle$
for some section $z$. The graded piece $R_{6}$ contains $xz$ and the four monomials of degree 6 in $x,y$. Assume by contradiction that there is a relation among these elements, so that $xf_{5}(x,y,z)=\lambda y^{3}$ for some $\lambda\in{\mathbb{C}}$. If $\lambda=0$, then $f_{5}=0$ by fact (a) and the linear relation is trivial. If $\lambda\neq 0$, then squaring the relation we obtain that $x^{2}$ divides $y^{6}$, but this is impossible since $y$ and $x^{2}$ have no common zero on $C$. So we see that we have no new generators in degree $d=6$, hence by Lemma 5.6, we have found all generators of $R$.
Let $u\in R_{6}$ be as in Proposition 3.3. We can write $u=xw+\lambda y^{3}$ with $w\in R_{5}$ and $\lambda\in{\mathbb{C}}$. Since $u\notin{\mathbb{C}}[x,y]$ as in the proof of Proposition 3.3 we have $w=\alpha z+p(x,y)$, with $\alpha\in{\mathbb{C}}^{*}$, so up to changing $z$ we may assume $u=xz+\lambda y^{3}$.
Now we study the relations. By Proposition 3.3, we may write $z^{2}=f_{10}(x^{2},y)+uf_{4}(x^{2},y)\in R_{10}$. Combining this with the relation $u^{2}-f_{12}(x^{2},y)=0$ in the canonical ring we have
$x^{2}f_{10}(x^{2},y)+ux^{2}f_{4}(x^{2},y)=x^{2}z^{2}=\lambda^{2}y^{6}-2\lambda y^{3}u+u^{2}=\lambda^{2}y^{6}-2\lambda uy^{3}+f_{12}(x^{2},y).$
A basis of in $R_{12}$ is $x^{12},x^{10}y,\dots,y^{6},x^{6}u,x^{4}yu,x^{2}y^{2}u,y^{3}u$, so comparing the coefficients of the monomials involving $u$ we see that $\lambda=0$, $f_{4}=0$, $xz=u$ and we have the claimed relation $z^{2}-f_{10}=0$. Comparing the Hilbert series (cf. Proposition 5.5), we see that we have found the full ring.
$\square$
5.C.2. Type $A(0)$
Now we prove that if $C$ is of type $A(0)$ then
$R\cong{\mathbb{C}}[y_{1},y_{2},z_{1},z_{2}]/(z_{1}^{2}-f_{6},z_{2}^{2}-g_{6}),$
with $\deg(y_{1},y_{2},z_{1},z_{2})=(2,2,3,3)$.
Proof.
Since $R_{1}=H^{0}({\mathcal{L}})=0$, we start with the basis $y_{1},y_{2}$ of $R_{2}=H^{0}(\omega_{C})$ coming from the canonical ring of $C$. We read off $\dim R_{3}=h^{0}({\mathcal{L}}(K_{C}))=2$ from the Hilbert series of $R$ (cf. Proposition 5.5) and choose a basis $z_{1},z_{2}$. Since ${\mathbb{C}}[y_{1},y_{2}]$ is a subring of $R$, we have ${\mathbb{C}}[y_{1},y_{2},z_{1},z_{2}]_{\leq 4}=R_{\leq 4}$. Since $h^{0}({\mathcal{L}})=h^{1}({\mathcal{L}})=0$ and $|K_{C}|$ is a free pencil, the multiplication $R_{2}\otimes R_{3}\to R_{5}$ is a bijection by Proposition 2.6, so
$R_{5}=\langle y_{1}z_{1},y_{2}z_{1},y_{1}z_{2},y_{2}z_{2}\rangle.$
From the canonical ring, $R_{6}$ is generated by cubic polynomials in $y_{1},y_{2}$ and by $u$, where $u$ is as in Proposition 3.3. Hence by Lemma 5.6 we have found all generators.
Thus the elements of $\operatorname{\mathsf{Sym}}^{2}(z_{1},z_{2})$ can be expressed in terms of $y_{1},y_{2},u$, namely there are three relations of the form
$z_{1}^{2}=\alpha u+a_{6}(y_{1},y_{2}),\quad z_{1}z_{2}=\beta u+b_{6}(y_{1},y_{2}),\quad z_{2}^{2}=\gamma u+c_{6}(y_{1},y_{2})$
Comparing the Hilbert series (cf. Proposition 5.5) we see that there are no additional relations. Assume that $\alpha=\beta=\gamma=0$: then $C$ contains the point $q=[0,0,0,0,1]$ and for a general choice of $a_{6},b_{6},c_{6}$ the tangent cone to $C$ at this point has six irreducible components, so $C$ does not have an $A_{n}$ singularity at $q$, contradicting Corollary 3.9. Therefore at least one of $\alpha,\beta,\gamma$ is nonzero and we can use one of the relations above to eliminate $u$. Up to a linear change of the coordinates $z_{1},z_{2}$ the remaining relations can be written either as:
(5.7) $z_{1}^{2}=f_{6}(y_{1},y_{2}),\quad z_{2}^{2}=g_{6}(y_{1},y_{2})$
or
$z_{1}^{2}=f_{6}(y_{1},y_{2}),\quad z_{1}z_{2}=h_{6}(y_{1},y_{2}).$
In the latter case, $C$ contains the point $[0,0,0,1]$ and again one can show that the corresponding singularity is not of type $A_{n}$. So the relations are as in (5.7). $\square$
Remark 5.8
If $({\mathcal{L}},\mu)$ is a ggs structure on a curve $C$ of type $A$ then it is possible to extend the action of the bicanonical involution $\iota$ on the canonical ring (see Corollary 3.6), to the half-canonical ring $R(C,\{{\mathcal{L}},\omega_{C}\})$. In case $A(1)$ we take
$(x,y,z)\mapsto(x,y,-z)$
and in case $A(0)$ we take
$(y_{1},y_{2},z_{1},z_{2})\mapsto(y_{1},y_{2},z_{1},-z_{2}).$
This amounts to choosing a linearisation of ${\mathcal{L}}$ compatible with the one chosen for $\omega_{C}$ (cf. Remark 3.7). Note that these are not the only possible choices: in the former case, we can also consider $(x,y,z)\mapsto(-x,y,-z)$ and in the latter case $(y_{1},y_{2},z_{1},z_{2})\mapsto(y_{1},y_{2},-z_{1},z_{2})$.
5.D. Type B: talking ’bout an involution
Now let $C$ be a canonical curve of type B, i.e., $C=C_{1}\cup C_{2}$ is a reducible and reduced Gorenstein curve of arithmetic genus two such that the two components have arithmetic genus one and intersect in a node $p$ (see Corollary  3.9). By Corollary 3.6, the bicanonical map $\varphi\colon C\to Q=Q_{1}\cup Q_{2}$ has degree $2$ with $Q_{i}\cong{\mathbb{P}}^{1}$, and the bicanonical involution $\iota$ is regular. We say a ggs structure $({\mathcal{L}},\mu)$ is of type $B(i)$ if $C$ is of type $B$ and $h^{0}(C,{\mathcal{L}})=i$. By Proposition 4.7 we have $i=0,1,2$.
Recall that we have chosen a linearisation on $\omega_{C}$ with respect to the bicanonical involution $\iota$ (Remark 3.7). In case $A$ we have determined all the possible half-canonical rings and a posteriori we have checked that the action of $\iota$ on the canonical ring extends to the half canonical rings. In other words, we have shown that every ggs structure $({\mathcal{L}},\mu)$ on a curve of type $A$ is linearisable and that the map $\mu\colon{\mathcal{L}}\otimes{\mathcal{L}}\to\omega_{C}$ is equivariant with respect to the fixed linearization of $\omega_{C}$. In case $B$ we proceed in a different order: we first show the existence of a linearisation such that the map $\mu\colon{\mathcal{L}}\otimes{\mathcal{L}}\to\omega_{C}$ is equivariant and then we use the induced action of $\iota$ on $R(C,\{{\mathcal{L}},\omega_{C}\})$ to analyse the ring structure.
Lemma 5.9 —
Let $C$ be a Gorenstein curve of genus two with $\omega_{C}$ ample of type $B$ and let $({\mathcal{L}},\mu)$ be a ggs structure on $C$. Then:
  1. (i)
    ${\mathcal{L}}$ can be linearised with respect to the bicanonical involution $\iota$ of $C$
  2. (ii)
    for any choice of linearisation of ${\mathcal{L}}$ the map $\mu\colon{\mathcal{L}}\otimes{\mathcal{L}}\to\omega_{C}$ is equivariant with respect to the linearisation of $\omega_{C}$ fixed in Remark 3.7.
Proof.
By Proposition 4.3 there are a partial normalisation $f\colon\widetilde{C}\to C$ and a regular ggs structure (~$(\widetilde{\mathcal{L}},\widetilde{\mu})$ on $\widetilde{C}$ such that ${\mathcal{L}}=f_{*}\widetilde{\mathcal{L}}$.
$(i)$ In the proof of Proposition 4.7, $(i)$ we have seen that $\iota$ induces an involution $\widetilde{\iota}$ of $\widetilde{C}$ such that $\widetilde{\iota}^{*}\widetilde{\mathcal{L}}\cong\widetilde{\mathcal{L}}$. Since $\widetilde{\mathcal{L}}$ is a line bundle, this implies that $\widetilde{\mathcal{L}}$ admits a linearisation. Pushing forward to $C$ we obtain a linearisation of ${\mathcal{L}}$.
$(ii)$ Linearisations of ${\mathcal{L}}$, $\widetilde{\mathcal{L}}$ with respect to $\iota$, $\widetilde{\iota}$ can be seen to be in bijection arguing as in the proof of Lemma 2.3, $(ii)$. By relative duality there is a ${\mathcal{O}}_{C}$-module homomorphism $f_{*}\omega_{\widetilde{C}}\to\omega_{C}$ which is invariant with respect to the natural linearisations of these sheaves with respect to $\widetilde{\iota}$ and $\iota$ (cf. Remark 3.7). So it is enough to show that for any given linearisation of $\widetilde{\mathcal{L}}$, the isomorphism ~~$\widetilde{\mathcal{L}}\otimes\widetilde{\mathcal{L}}\to\omega_{\widetilde{C}}$ induces on $\omega_{\widetilde{C}}$ the opposite of the natural linearisation.
By [FFP22, Lemma 4.4] there is a trivializing open cover $\{U_{j}\}$ for $\widetilde{\mathcal{L}}$ on $\widetilde{C}$ such that the $U_{j}$ are $\widetilde{\iota}$-invariant. Let $\sigma_{j}$ be a local generator of ~$\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{U_{j}}$}}$ and let $q\in U_{j}$ be a point: then at least one between $\sigma_{j}+{\widetilde{\iota}}^{*}\sigma_{j}$ and $\sigma_{j}-{\widetilde{\iota}}^{*}\sigma_{j}$ does not vanish at $q$. So, up to refining the cover, we may assume in addition that for every $j$ the sheaf ~$\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{U_{j}}$}}$ is generated by an element $\tau_{j}$ such that ${\widetilde{\iota}}^{*}\tau_{j}=\pm\tau_{j}$. Via the isomorphism ~~$\widetilde{\mathcal{L}}\operatorname{\otimes}\widetilde{\mathcal{L}}\to\omega_{\widetilde{C}}$ we obtain a local trivialization of $\omega_{\widetilde{C}}$ such that for every $j$ there is a local generator $\alpha_{j}$ on $U_{j}$ that verifies ${\widetilde{\iota}}^{*}\alpha_{j}=\alpha_{j}$. In other words, $\omega_{\widetilde{C}}$ is the pull back of a line bundle $\widetilde{\mathcal{M}}$ on $\widetilde{C}/\widetilde{\iota}$ and the linearisation induced by the ggs structure coincides with the linearisation as the pullback of $\widetilde{\mathcal{M}}$. By Proposition 4.5 ${\mathcal{L}}$ is not regular at the intersection point of $C_{1}$ and $C_{2}$, hence $\widetilde{C}=\widetilde{C}_{1}\sqcup\widetilde{C}_{2}$ is a disjoint union of curves of arithmetic genus 0 or 1. The bicanonical involution $\iota$ induces an involution $\widetilde{\iota}$ of $\widetilde{C}$ that preserves the two components and such that $Q_{i}:=\widetilde{C}_{i}/\widetilde{\iota}$ is smooth rational for $i=1,2$. So the quotient map $\varphi_{i}\colon\widetilde{C}_{i}\to Q_{i}$ is flat and ${\varphi_{i}}_{*}{\mathcal{O}}_{\widetilde{C}_{i}}={\mathcal{O}}_{Q_{i}}\oplus{\mathcal{O}}_{Q_{i}}(-a)$, where $a=p_{a}(C_{i})+1$. Relative duality gives ${\varphi_{i}}_{*}\omega_{\widetilde{C}_{i}}=\omega_{Q_{i}}\oplus\omega_{Q_{i}}(a)$. Since $\deg\omega_{\widetilde{C}_{i}}=2\deg{\mathcal{M}}{\raisebox{-0.1pt}{$|_{Q_{i}}$}}$, we conclude that ${\mathcal{M}}{\raisebox{-0.1pt}{$|_{Q_{i}}$}}=\omega_{Q_{i}}(a)$, namely the linearisation induced on $\omega_{C_{i}}$ by the isomorphism ~~$\widetilde{\mathcal{L}}\operatorname{\otimes}\widetilde{\mathcal{L}}\to\omega_{\widetilde{C}}$ is the opposite of the natural one. $\square$
Lemma 5.10 —
Let $({\mathcal{L}},\mu)$ be a ggs structure of type $B(i)$ with $i=h^{0}({\mathcal{L}})=a_{1}+a_{2}$ where $(a_{1},a_{2})=(0,0)$, $(0,1)$ or $(1,1)$. Then, if we use the linearisation of $\omega_{C}$ chosen in Remark 3.7:
  1. (i)
    there is a unique linearisation of ${\mathcal{L}}$ such that both $R_{3}=H^{0}({\mathcal{L}}(\omega_{C}))$ and $R_{1}=H^{0}({\mathcal{L}})$ are $\iota$-invariant;
  2. (ii)
    for the above choice of linearisation we have the following splittings under the action of $\iota$:
If $n>0$ is odd:
$R_{2n+1}^{+}$ $\cong H^{0}\left({\mathcal{O}}_{Q_{1}}{\left(\tfrac{n-1}{2}\right)}\right)\oplus H^{0}\left({\mathcal{O}}_{Q_{2}}{\left(\tfrac{n-1}{2}\right)}\right)$
$R_{2n+1}^{-}$ $\cong H^{0}\left({\mathcal{O}}_{Q_{1}}{\left(\tfrac{n-1}{2}-1\right)}\right)\oplus H^{0}\left({\mathcal{O}}_{Q_{2}}{\left(\tfrac{n-1}{2}-1\right)}\right)$
If $n\geq 0$ is even:
$R_{2n+1}^{+}$ $\cong H^{0}\left({\mathcal{O}}_{Q_{1}}{\left(\tfrac{n}{2}-1+a_{1}\right)}\right)\oplus H^{0}\left({\mathcal{O}}_{Q_{2}}{\left(\tfrac{n}{2}-1+a_{2}\right)}\right)$
$R_{2n+1}^{-}$ $\cong H^{0}\left({\mathcal{O}}_{Q_{1}}{\left(\tfrac{n}{2}-1-a_{1}\right)}\right)\oplus H^{0}\left({\mathcal{O}}_{Q_{2}}{\left(\tfrac{n}{2}-1-a_{2}\right)}\right)$
Proof.
Since ${\mathcal{L}}$ is not locally free at the intersection point $p$ of $C_{1}$ and $C_{2}$ (cf. Proposition 4.5), by Proposition 4.3 there is a partial normalisation $f\colon\widetilde{C}_{1}\sqcup\widetilde{C}_{2}\to C$ such that ${\mathcal{L}}=f_{*}\widetilde{\mathcal{L}}$, where ~=~$\widetilde{\mathcal{L}}=\widetilde{\mathcal{L}}_{1}\oplus\widetilde{\mathcal{L}}_{2}$ and ~$\widetilde{\mathcal{L}}_{i}$ is a line bundle supported on $\widetilde{C}_{i}$. Recall that $\widetilde{C}_{i}$ is Gorenstein, $p_{a}(\widetilde{C}_{i})=0$ or 1 and ~$\widetilde{\mathcal{L}}_{i}^{\otimes 2}=\omega_{\widetilde{C}_{i}}$. We set ${\mathcal{L}}_{i}:=f_{*}\widetilde{\mathcal{L}}_{i}$ and $a_{i}:=h^{0}({\mathcal{L}}_{i})=h^{0}(\widetilde{\mathcal{L}}_{i})$. We have $a_{1}+a_{2}=h^{0}({\mathcal{L}})\leq 2$ by Proposition 4.7, so if we assume $a_{1}\leq a_{2}$ the possibilities for the pair $(a_{1},a_{2})$ are $(0,0)$, $(0,1)$ or $(1,1)$.
As we have seen in the proof of Proposition 4.7, $\iota$ induces an involution $\widetilde{\iota}$ on $\widetilde{C}_{1}\sqcup\widetilde{C}_{2}$ preserving the two components and with $\widetilde{C}_{i}/\widetilde{\iota}=Q_{i}\cong{\mathbb{P}}^{1}$. If $\varphi_{i}\colon\widetilde{C}_{i}\to Q_{i}$ is the quotient map, i.e., $\varphi_{i}$ is the restriction to $\widetilde{C}_{i}$ of the composition $\varphi\circ f\colon\widetilde{C}_{1}\sqcup\widetilde{C}_{2}\to Q_{1}\cup Q_{2}$, then we have
$\omega_{C}^{\otimes n}\otimes{\mathcal{L}}=\omega_{C}^{\otimes n}\otimes({\mathcal{L}}_{1}\oplus{\mathcal{L}}_{2})=(\omega_{C}^{\otimes n}\otimes{\mathcal{L}}_{1})\oplus(\omega_{C}^{\otimes n}\oplus{\mathcal{L}}_{2})=$
$(\omega_{C}^{\otimes n}{\raisebox{-0.1pt}{$|_{C_{1}}$}}\otimes{\mathcal{L}}_{1})\oplus(\omega_{C}^{\otimes n}{\raisebox{-0.1pt}{$|_{C_{2}}$}}\otimes{\mathcal{L}}_{2})={\mathcal{L}}_{1}\otimes{\mathcal{O}}_{C_{1}}(np)\oplus{\mathcal{L}}_{2}\otimes{\mathcal{O}}_{C_{2}}(np).$
Therefore
$R_{2n+1}=H^{0}(\omega_{C}^{\otimes n}\otimes{\mathcal{L}})\cong H^{0}\left({\varphi_{1}}_{*}{\widetilde{\mathcal{L}}_{1}(np_{1})}\right)\oplus H^{0}\left({\varphi_{2}}_{*}{\widetilde{\mathcal{L}}_{2}(np_{2})}\right).$
where $p_{1},p_{2}$ are the preimages of the node $p=C_{1}\cap C_{2}$, In turn, each summand $H^{0}\left({\varphi_{i}}_{*}{\widetilde{\mathcal{L}}_{i}(np_{i})}\right)$ decomposes into eigenspaces under the action of $\iota$. Since ${\mathcal{O}}_{C_{i}}(2p_{i})=\varphi_{i}^{*}{\mathcal{O}}_{Q_{i}}(1)$, the projection formula gives
${\varphi_{i}}_{*}{{\mathcal{L}}_{i}(np_{i})}=\begin{cases}({\varphi_{i}}_{*}{\mathcal{L}}_{i})\otimes{\mathcal{O}}_{Q_{i}}(\tfrac{n}{2})&\text{ $n$ even,}\\ ({\varphi_{i}}_{*}{\mathcal{L}}_{i}(p_{i}))\otimes{\mathcal{O}}_{Q_{i}}(\tfrac{n-1}{2})&\text{ $n$ odd.}\end{cases}$
In addition, ${\mathcal{O}}_{C_{i}}(2p_{i})$ has the pull-back linearisation, so it is enough to determine the decomposition into eigensheaves of ${\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}$ and ${\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}(p_{i})$.
The line bundle ~$\widetilde{\mathcal{L}}_{i}(p_{i})$ has degree $p_{a}(\widetilde{C}_{i})$. If $p_{a}(\widetilde{C}_{i})=0$, then ~$\widetilde{\mathcal{L}}_{i}(p_{i})={\mathcal{O}}_{\widetilde{C}_{i}}$ and ${\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}(p_{i})={\mathcal{O}}_{Q_{i}}\oplus{\mathcal{O}}_{Q_{i}}(-1)$. If $p_{a}(\widetilde{C}_{i})=1$, then $h^{0}(\widetilde{\mathcal{L}}_{i}(p_{i}))=1$ and $h^{1}(\widetilde{\mathcal{L}}_{i}(p_{i}))=0$, so that ${\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}(p_{i})={\mathcal{O}}_{Q_{i}}\oplus{\mathcal{O}}_{Q_{i}}(-1)$ also in this case. So for $i=1,2$ we choose the linearization of ~$\widetilde{\mathcal{L}}_{i}(p_{i})$ in such a way that ${\mathcal{O}}_{Q_{i}}$ is the invariant summand; this is the same as choosing a linearisation of ${\mathcal{L}}\otimes\omega_{C}$ as claimed in $(i)$. Note that choosing a linearisation of ${\mathcal{L}}\otimes\omega_{C}$ is the same as choosing a linearisation of ${\mathcal{L}}$, because $\omega_{C}$ is invertible and has a chosen linearisation (cf. Remark 3.7 and Lemma 5.9). Note also that if $H^{0}({\mathcal{L}})$ contains a nonzero section $x$, then $x^{3}\in H^{0}({\mathcal{L}}\otimes\omega_{C})$ is invariant, hence so is $x$ and claim (i) is proven. Statement $(ii)$ for odd $n>0$ follows from the above remarks.
We consider now ${\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}$: we have $a_{i}=h^{0}(\widetilde{\mathcal{L}}_{i})=h^{1}(\widetilde{\mathcal{L}}_{i})$, so we have
${\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}={\mathcal{O}}_{Q_{i}}(-1+a_{i})\oplus{\mathcal{O}}_{Q_{i}}(-1-a_{i}).$
If $a_{i}=0$, the two eigensheaves are isomorphic and claim $(ii)$ for even $n$ follows immediately. If $a_{i}=1$, by $(i)$ we know that the invariant subsheaf is the first one and we get $(ii)$ also in this case. $\square$
We are now going to prove cases $B(0),B(1)$ and $B(2)$ of Theorem 5.2.
5.D.1. Type $B(1)$
Let $({\mathcal{L}},\mu)$ be a ggs curve of type $B(1)$ so that $(a_{1},a_{2})=(0,1)$. Here we prove that $R(C,\{{\mathcal{L}},\omega_{C}\})$ is isomorphic to ${\mathbb{C}}[x,y,w,v,z,u]/I$ where the degrees are $\deg(x,y,w,v,z,u)=(1,2,3,4,5,6)$ and $I$ is generated by the following relations:
2×2 minors of ($2\times 2\text{ minors of }\begin{pmatrix}0&y&w&z\\ x&w&v&u\end{pmatrix}\ \text{ and }\begin{array}[]{rcl}z^{2}&=&yg_{8}(y,v)\\ zu&=&wg_{8}(y,v)\\ u^{2}&=&vg_{8}(y,v)+x^{4}h_{8}(x,v)\end{array}$
With respect to the linearisation of Lemma 5.10 the involution $\iota$ acts on $R(C,\{{\mathcal{L}},\omega_{C}\})$ by $(x,y,w,v,z,u)\mapsto(x,y,w,v,-z,-u)$.
Let us denote the non-zero section of ${\mathcal{L}}$ by $x$. By Lemma 5.10 the section $x$ is invariant and supported on $C_{2}$.
Since $C$ is reduced we have $x^{2}\neq 0$ so that by Proposition 3.3 and its proof we can write
$R(C,K_{C})\cong{\mathbb{C}}[x^{2},y,v,u]/(x^{2}y,u^{2}-f_{12}(x^{2},y,v))\supset R(C,K_{C})^{+}={\mathbb{C}}[x^{2},y,v]/(x^{2}y).$
Note that $xy=0$, because $x$ and $y$ are supported on different components of $C$.
Now we have to add the remaining elements of odd degree. By Lemma 5.10 $R_{3}$ is invariant, generated by $x^{3}$ and by an element $w$ supported on $C_{1}$. Again, since $x$ and $w$ are supported on different components of $C$, we get $xw=0$.
Note that $R_{4}=R_{4}^{+}=\langle x^{4},y^{2},v\rangle$ and that $v$ is non-zero at the generic point of both components of $C$, in particular $v$ is not a zero-divisor in $R$.
Let us now look in degree $5$, where we have the decompositions
$R_{5}^{+}\cong H^{0}({{\mathcal{O}}_{Q_{1}}})\oplus H^{0}({{\mathcal{O}}_{Q_{2}}}(1))\text{ and }R_{5}^{-}\cong H^{0}({\mathcal{O}}_{Q_{1}})\oplus H^{0}({\mathcal{O}}_{Q_{2}}(-1)).$
The first summand of $R_{5}^{+}$ is $\langle yw\rangle$, the second is $\langle x^{5},xv\rangle$ and we have a new anti-invariant generator $z$ of degree $5$ supported on $C_{1}$, which spans $R_{5}^{-}$.
We need the following Lemma:
Lemma 5.11 —
If $s,t\in R$ and $yt\neq 0$ then $s=0$ if and only if $xs=ts=0$.
Proof.
We are dealing with sections of torsion-free sheaves, so it is enough to check this at the generic point of each component. But $C$ has two (reduced) components and either $x$ or $y$ is non-zero at the generic point of each of them. $\square$
End of the proof of Theorem 5.2: Type $B(1)$.
We have analysed the situation up to degree $5$ already, so we start in degree $6$. By Proposition 3.3, $R_{6}$ is generated by $x^{2},y,v,u$ and $u$ does not vanish at the generic points of $C_{1}$ and $C_{2}$. Clearly, $w^{2}$ is an invariant section which is supported on $C_{1}$, so we get a relation
(5.12) $w^{2}=\alpha_{1}y^{3}+\alpha_{2}yv$
Moreover, $x$ and $z$ are supported on different components of $C$ so $xz=0$. We have accounted for all elements in $R_{6}$.
Using Lemma 5.10, we claim that $R_{7}$ decomposes as follows
$R_{7}^{+}=\langle wy^{2},wv\rangle\oplus\langle x^{7},x^{3}v\rangle\text{ and }R_{7}^{-}=\langle yz\rangle\oplus\langle xu\rangle,$
thus there are no new relations in degree $7$. Indeed, consider the first summand of $R_{7}^{+}$ and assume that there is a linear relation $\alpha wy^{2}+\beta wv=0$. At the generic point of $C_{1}$ we can divide by $w$ to get $\alpha y^{2}+\beta v=0$ on $C_{1}$, namely $\alpha y^{2}+\beta v=\gamma x^{4}$ for some scalar $\gamma$. Since $y^{2},x^{4},v$ are a basis of $R_{4}$ we get $\alpha=\beta=\gamma=0$ and $wy^{2},wv$ are independent. We can argue in a similar way for the second summand of $R_{7}$, dividing by $x^{3}$. The two summands of $R_{7}^{-}$ are easily treated by noting that $xu\neq 0$ and $yz\neq 0$.
We derive the following decomposition in degree $8$ from the canonical ring:
$R_{8}^{+}=\langle x^{8},x^{4}v,v^{2},y^{4},y^{2}v\rangle\text{ and }R_{8}^{-}=\langle x^{2}u,yu\rangle.$
The element $wz\in R_{8}^{-}$ is non-zero since both $w$ and $z$ are non-zero at the generic point of $C_{1}$, hence up to rescaling $w$ we can write
(5.13) $wz=yu.$
We skip over degree $9$ for now. Again, from our knowledge of the canonical ring, we know that there are no new generators in degree $10$. The only element not accounted for so far is $z^{2}$, which is non-zero, invariant, and supported on $C_{1}$, thus
(5.14) $z^{2}=\mu_{1}y^{5}+\mu_{2}y^{3}v+\mu_{3}yv^{2}.$
We now analyse the interaction between the relations found so far in order to clean up some coefficients. Taking the square of (5.13) and substituting (5.12), (5.14) and the relation $u^{2}=f_{12}$ in the canonical ring we get
(5.15) $y^{2}f_{12}=(yu)^{2}=w^{2}z^{2}=(\alpha_{1}y^{3}+\alpha_{2}yv)(\mu_{1}y^{5}+\mu_{2}y^{3}v+\mu_{3}yv^{2}).$
Since $f_{12}$ contains the monomial $v^{3}$ with non-zero coefficient by Proposition 3.3, we have $\alpha_{2}\mu_{3}\neq 0$. With the coordinate transformation $(\alpha_{1}y^{2}+\alpha_{2}v)\mapsto v$ the relation (5.12) becomes
(5.16) $w^{2}=yv.$
Substituting back into (5.15) we get
$y^{2}f_{12}=yv(\mu_{1}y^{5}+\mu_{2}y^{3}v+\mu_{3}yv^{2})=y^{2}vg_{8}(y,v)$
and therefore
$f_{12}=vg_{8}(y,v)+x^{4}h_{8}(x,v),\quad z^{2}=yg_{8}(y,v).$
Note that $y^{6}$ has been eliminated from the equation by our choice of $v$.
Now multiplying (5.13) by $z$ and substituting (5.14), we get
$y(uz)=w(z^{2})=ywg_{8}$
and since $uz$, $w$ and $y$ are supported on $C_{1}$, we may divide by $y$ to obtain
$uz=wg_{8}.$
Note that by Lemma 5.6, we have now found all generators of $R$ with the possible exception of degree $9$, which we now check as promised. We claim that the decomposition of Lemma 5.10 is:
$R_{9}^{+}=\langle wy^{3},wyv\rangle\oplus\langle x^{9},x^{5}v,xv^{2}\rangle,\quad R_{9}^{-}=\langle y^{2}z,vz\rangle\oplus\langle x^{3}u\rangle.$
We proceed just as we did in degree $7$. At the generic point of $C_{1}$, we can divide by $w$ to show that $wyv$ and $wy^{3}$ are linearly independent. Hence the first summand of $R_{9}^{+}$ is as claimed. For the second summand, we can divide by $x$ at the generic point of $C_{2}$.
None of the given elements of $R_{9}^{-}$ are zero by Lemma 5.11. Moreover, we may divide by $z$ to get linear independence of $y^{2}z,vz$ at the generic point of $C_{1}$. Thus $R_{9}$ decomposes as claimed.
Now $wu$ is also an element of $R_{9}$: it is non-zero, anti-invariant and supported on $C_{1}$, hence there is a final relation $wu=\lambda_{1}y^{2}z+\lambda_{2}vz$. Multiplying by $y$ and substituting (5.13) and (5.16) we get $yvz=ywu=y(\lambda_{1}y^{2}z+\lambda_{2}vz)=\lambda_{1}y^{3}z+\lambda_{2}yvz$ hence $\lambda_{1}=0$, $\lambda_{2}=1$ and the relation is
$vz=wu.$
Clearly, the given relations are contained in the kernel of the map, so it remains to show that there are no others. To do this, we compute the Hilbert series of ${\mathbb{C}}[x,y,w,v,z,u]/I$ (for example, using a computer algebra package), and check that it matches that of $R$.
5.D.2. Type $B(0)$
Let $({\mathcal{L}},\mu)$ be a ggs structure of type $B(0)$, so that $(a_{1},a_{2})=(0,0)$. We will prove that $R\cong{\mathbb{C}}[y_{1},y_{2},z_{1},z_{2},v,t_{1},t_{2},u]/I$ where the degrees are $\deg(y_{1},y_{2},z_{1},z_{2},v,t_{1},t_{2},u)=(2,2,3,3,4,5,5,6)$ and $I$ is generated by the $2\times 2$ minors of the symmetric $4\times 4$ matrix
$\begin{pmatrix}y_{1}&0&z_{1}&t_{1}\\ 0&y_{2}&z_{2}&t_{2}\\ z_{1}&z_{2}&v&u\\ t_{1}&t_{2}&u&g_{8}\end{pmatrix}$
With respect to the linearisation of Lemma 5.10 the involution $\iota$ acts on $R(C,\{{\mathcal{L}},\omega_{C}\})$ by $(y_{1},y_{2},z_{1},z_{2},v,t_{1},t_{2},u)\mapsto(y_{1},y_{2},z_{1},z_{2},v,-t_{1},-t_{2},-u)$.
Recall from Proposition 3.3 and Corollary 3.6 that the canonical ring of $C$ and its invariant subring are
$R(C,K_{C})\cong{\mathbb{C}}[y_{1},y_{2},v,u]/(y_{1}y_{2},u^{2}-f_{12}(y_{1},y_{2},v))\supset R(C,K_{C})^{+}={\mathbb{C}}[y_{1},y_{2},v]/(y_{1}y_{2}).$
We also need the following Lemma, whose proof is similar to that of Lemma 5.11.
Lemma 5.17 —
If $s,t\in R$ and $y_{1}t\neq 0$ then $s=0$ if and only if $y_{2}s=ts=0$.
End of the proof of Theorem 5.2: Type $B(0)$.
Clearly $R_{1}=0$ and $R_{2}=R_{2}^{+}=\langle y_{1},y_{2}\rangle$. The decomposition for $R_{3}$ given in Lemma 5.10 yields
$R_{3}^{+}\cong H^{0}({{\mathcal{O}}_{Q_{1}}})\oplus H^{0}({\mathcal{O}}_{Q_{2}})\text{ and }R_{3}^{-}\cong H^{0}({{\mathcal{O}}_{Q_{1}}}(-1))\oplus H^{0}({\mathcal{O}}_{Q_{2}}(-1)),$
so we choose generators $z_{1}$, $z_{2}$ for $R_{3}=R_{3}^{+}$. These are supported on different components of $C$, so $z_{1}z_{2}=0$. Similarly we get relations $y_{1}z_{2}=y_{2}z_{1}=0$.
Note that $R_{4}=R_{4}^{+}=\langle y_{1}^{2},y_{2}^{2},v\rangle$ and that $v$ is non-zero at the generic point of both components of $C$, in particular $v$ is not a zero-divisor in $R$.
In degree $5$, we have
$R_{5}^{+}$ $\cong H^{0}({{\mathcal{O}}_{Q_{1}}})\oplus H^{0}({{\mathcal{O}}_{Q_{2}}})=\langle y_{1}z_{1}\rangle\oplus\langle y_{2}z_{2}\rangle$
$R_{5}^{-}$ $\cong H^{0}({\mathcal{O}}_{Q_{1}})\oplus H^{0}({\mathcal{O}}_{Q_{2}})=\langle t_{1}\rangle\oplus\langle t_{2}\rangle$
where the new generators $t_{1},t_{2}$ of degree $5$ are anti-invariant because $y_{i}z_{i}$ is invariant for $i=1,2$. The $t_{i}$ are supported on different components of $C$, so we again have relations $t_{1}t_{2}=0$ and $t_{i}z_{j}=t_{i}y_{j}=0$ for $i\neq j$.
In degree $6$ we know from the canonical ring that
$R_{6}^{+}=\langle y_{1}^{3},y_{2}^{3},y_{1}v,y_{2}v\rangle\text{ and }R_{6}^{-}=\langle u\rangle.$
Since $z_{1}^{2}$, $z_{2}^{2}$ are in $R_{6}^{+}$ and supported on $C_{1}$ respectively $C_{2}$, we have two relations
(5.18) $z_{1}^{2}=\alpha_{1}y_{1}^{3}+\alpha_{2}y_{1}v,\quad\quad z_{2}^{2}=\beta_{1}y_{2}^{3}+\beta_{2}y_{2}v.$
In the usual way, $R_{7}$ decomposes as
$R_{7}^{+}$ $\cong H^{0}({\mathcal{O}}_{Q_{1}}(1))\oplus H^{0}({\mathcal{O}}_{Q_{2}}(1))=\langle y_{1}^{2}z_{1},z_{1}v\rangle\oplus\langle y_{2}^{2}z_{2},z_{2}v\rangle$
$R_{7}^{-}$ $\cong H^{0}({\mathcal{O}}_{Q_{1}})\oplus H^{0}({\mathcal{O}}_{Q_{2}})=\langle y_{1}t_{1}\rangle\oplus\langle y_{2}t_{2}\rangle$
Indeed $y_{i}t_{i}\neq 0$ and $y_{1}^{2}z_{1},z_{1}v$ are linearly independent because we can divide by $z_{1}$ at the generic point of $C_{1}$ to get $y_{1}^{2}$, $v$ which are linearly independent on $C_{1}$.
From the involution on the canonical ring, we have
$R_{8}^{+}=\langle y_{1}^{4},y_{2}^{4},y_{1}^{2}v,y_{2}^{2}v,v^{2}\rangle\ \text{ and }\ R_{8}^{-}=\langle y_{1}u,y_{2}u\rangle.$
Now $z_{i}t_{i}$ are non-zero elements of $R_{8}$ which are respectively supported on $C_{i}$ for $i=1,2$. Since $z_{i}t_{i}$ is anti-invariant for $i=1,2$, we have relations
(5.19) $z_{1}t_{1}=y_{1}u,\quad\quad z_{2}t_{2}=y_{2}u$
where we have rescaled $y_{1},y_{2}$ to tidy up the coefficients.
Next we consider degree $9$ where we have the decomposition
$R_{9}^{+}$ $\cong H^{0}({\mathcal{O}}_{Q_{1}}(1))\oplus H^{0}({\mathcal{O}}_{Q_{2}}(1))=\langle y_{1}^{3}z_{1},y_{1}z_{1}v\rangle\oplus\langle y_{2}^{3}z_{2},y_{2}z_{2}v\rangle,$
$R_{9}^{-}$ $\cong H^{0}({\mathcal{O}}_{Q_{1}}(1))\oplus H^{0}({\mathcal{O}}_{Q_{2}}(1))=\langle y_{1}^{2}t_{1},vt_{1}\rangle\oplus\langle y_{2}^{2}t_{2},vt_{2}\rangle.$
As usual, we can show that the listed monomials are linearly independent by dividing by some common factor at the generic point of $C_{i}$. Now for $i=1,2$ we know that $z_{i}u$ in $R_{9}$ is non-zero, anti-invariant and is supported on $C_{i}$. Hence $z_{i}u$ is in the summand $\langle y_{i}^{2}t_{i},vt_{i}\rangle$. Thus we have relations
(5.20) $z_{1}u=\gamma_{1}y_{1}^{2}t_{1}+\gamma_{2}vt_{1},\quad\quad z_{2}u=\delta_{1}y_{2}^{2}t_{2}+\delta_{2}vt_{2}$
Since there are no new generators in degree $9$, it follows from Lemma 5.6, that $R$ is generated in degrees $\leq 6$.
We have $t_{1}^{2},t_{2}^{2}$ in $R_{10}^{+}$ each supported on $C_{i}$ for $i=1,2$. Thus from the canonical ring we get relations
(5.21) $t_{1}^{2}=\lambda_{1}y_{1}^{5}+\lambda_{2}y_{1}^{3}v+\lambda_{3}y_{1}v^{2},\quad\quad t_{2}^{2}=\mu_{1}y_{2}^{5}+\mu_{2}y_{2}^{3}v+\mu_{3}y_{2}v^{2}$
Now we study the interactions between relations and clean up several coefficients. Taking the squares of (5.19) and substituting (5.18), (5.21) and the relation $u^{2}=f_{12}$ from the canonical ring, we get the relations
(5.22) $\begin{split}y_{1}^{2}f_{12}&=y_{1}^{2}u^{2}=z_{1}^{2}t_{1}^{2}=(\alpha_{1}y_{1}^{3}+\alpha_{2}y_{1}v)(\lambda_{1}y_{1}^{5}+\lambda_{2}y_{1}^{3}v+\lambda_{3}y_{1}v^{2})\\ y_{2}^{2}f_{12}&=y_{2}^{2}u^{2}=z_{2}^{2}t_{2}^{2}=(\beta_{1}y_{2}^{3}+\beta_{2}y_{2}v)(\mu_{1}y_{2}^{5}+\mu_{2}y_{2}^{3}v+\mu_{3}y_{2}v^{2})\end{split}$
Since $f_{12}$ contains the monomial $v^{3}$ with non-zero coefficient it follows that $\alpha_{2}$ and $\beta_{2}$ must be non-zero. By a coordinate change of the form $v\mapsto v+\frac{\alpha_{1}}{\alpha_{2}}y_{1}^{2}+\frac{\beta_{1}}{\beta_{2}}y_{2}^{2}$, noting that $y_{1}y_{2}=0$ and by rescaling $z_{1},z_{2}$, the equations (5.18) become
(5.23) $z_{1}^{2}=y_{1}v,\quad\quad z_{2}^{2}=y_{2}v.$
Substituting back into (5.22) we get
(5.24) $\begin{split}y_{1}^{2}f_{12}&=y_{1}v(\lambda_{1}y_{1}^{5}+\lambda_{2}y_{1}^{3}v+\lambda_{3}y_{1}v^{2})=y_{1}^{2}v(\lambda_{1}y_{1}^{4}+\lambda_{2}y_{1}^{2}v+\lambda_{3}v^{2})\\ y_{2}^{2}f_{12}&=y_{2}v(\mu_{1}y_{2}^{5}+\mu_{2}y_{2}^{3}v+\mu_{3}y_{2}v^{2})=y_{2}^{2}v(\mu_{1}y_{2}^{4}+\mu_{2}y_{2}^{2}v+\mu_{3}v^{2})\end{split}$
Therefore
$f_{12}(y_{1},y_{2},v)=vg_{8}(y_{1},y_{2},v),\quad\quad u^{2}=vg_{8}.$
It follows immediately from the above definition of $g_{8}$ and the relation $y_{1}y_{2}=0$, that the equations (5.21) become
(5.25) $t_{1}^{2}=y_{1}g_{8},\quad\quad t_{2}^{2}=y_{2}g_{8}.$
Multiplying (5.25) by $u$ and substituting (5.19) gives
(5.26) $t_{1}^{2}u=y_{1}ug_{8}=z_{1}t_{1}g_{8},\quad\quad t_{2}^{2}u=y_{2}ug_{8}=z_{2}t_{2}g_{8}$
and dividing by $t_{i}$ at the generic point of $C_{j}$ yields relations
(5.27) $t_{1}u=z_{1}g_{8},\quad\quad t_{2}u=z_{2}g_{8}.$
Multiplying (5.20) by $y_{i}$ and substituting (5.19) and (5.23) we get
(5.28) $\begin{split}y_{1}vt_{1}&=z_{1}^{2}t_{1}=y_{1}z_{1}u=\gamma_{1}y_{1}^{3}t_{1}+\gamma_{2}y_{1}vt_{1}\\ y_{2}vt_{2}&=z_{2}^{2}t_{2}=y_{2}z_{2}u=\delta_{1}y_{2}^{3}t_{2}+\delta_{2}y_{2}vt_{2}\end{split}$
hence $\gamma_{1}=\delta_{1}=0$ and $\gamma_{2}=\delta_{2}=1$ and the relations are
$z_{1}u=vt_{1},\quad\quad z_{2}u=vt_{2}.$
To show that there are no further relations, we computed the Hilbert series of ${\mathbb{C}}[y_{1},y_{2},z_{1},z_{2},v,t_{1},t_{2},u]/(\text{relations})$ and checked that it matches with that of $R$.
5.D.3. Type $B(2)$
Let $({\mathcal{L}},\mu)$ be a ggs structure of type $B(2)$. We show that $R\cong{\mathbb{C}}[x_{1},x_{2},v,u]/\left(x_{1}x_{2}=0,u^{2}=f_{12}(x_{1},x_{2},v)\right)$. With respect to the linearisation of Lemma 5.10 the involution $\iota$ acts on $R(C,\{{\mathcal{L}},\omega_{C}\})$ by $(x_{1},x_{2},v,u)\mapsto(x_{1},x_{2},v,-u)$.
Proof.
We start with a basis for $R_{1}=\langle x_{1},x_{2}\rangle$, because $h^{0}({\mathcal{L}})=2$. Since $C$ has arithmetic genus $2$, we see that $x_{1}^{2},x_{1}x_{2},x_{2}^{2}$ are not linearly independent in $R_{2}$. The curve $C$ is reduced so we may choose coordinates such that $x_{1}x_{2}=0$ and $R_{2}=R_{2}^{+}=\langle x_{1}^{2},x_{2}^{2}\rangle$ where $x_{1}^{2}=y_{1}$ and $x_{2}^{2}=y_{2}$. We read off the dimension of $R_{n}$ using the Hilbert series and continue with $R_{3}=\langle x_{1}^{3},x_{2}^{3}\rangle$ and $R_{4}=R_{4}^{+}=\langle x_{1}^{4},x_{2}^{4},v\rangle$.
We tabulate the rest of the computation up to degree $9$, using the canonical ring and the decomposition of $R_{2n+1}$ from Lemma 5.10.
$d$ $R_{d}^{+}$ $R_{d}^{-}$
$5$ $\langle x_{1}^{5},x_{1}v\rangle\oplus\langle x_{2}^{5},x_{2}v\rangle$
$6$ $\langle x_{1}^{6},x_{1}^{2}v\rangle\oplus\langle x_{2}^{6},x_{2}^{2}v\rangle$ $\langle u\rangle$
$7$ $\langle x_{1}^{7},x_{1}^{3}v\rangle\oplus\langle x_{2}^{7},x_{2}^{3}v\rangle$ $\langle x_{1}u\rangle\oplus\langle x_{2}u\rangle$
$8$ $\langle x_{1}^{8},x_{1}^{4}v\rangle\oplus\langle x_{2}^{8},x_{2}^{4}v\rangle\oplus\langle v^{2}\rangle$ $\langle x_{1}^{2}u\rangle\oplus\langle x_{2}^{2}u\rangle$
$9$ $\langle x_{1}^{9},x_{1}^{5}v,x_{1}v^{2}\rangle\oplus\langle x_{2}^{9},x_{2}^{5}v,x_{2}v^{2}\rangle$ $\langle x_{1}^{3}u\rangle\oplus\langle x_{2}^{3}u\rangle$
The direct summands group together basis elements that are supported on either one of the curves respectively both of them. For each $R_{2m+1}$, the listed elements form a basis because we can divide by $x_{i}$ at the generic point of the component $C_{i}$ to get the basis of $R_{2m}$, which is a summand of the canonical ring. By Lemma 5.6, we have found all generators of $R$.
The relation $u^{2}=f_{12}(x_{1},x_{2},v)$ is induced by the relation in the canonical ring of type B (cf. Proposition 3.3 and Corollary 3.6).
Clearly, the Hilbert series of ${\mathbb{C}}[x_{1},x_{2},v,u]/\left(x_{1}x_{2}=0,u^{2}=f_{12}(x_{1},x_{2},v)\right)$ matches that of $R$, so the proof is finished. $\square$
5.E. The “converse” statement of Theorem 5.2
In this section we are going to show that each ideal given in Table 2 defines a reduced Gorenstein curve $C$ of genus two such that $\omega_{C}={\mathcal{O}}_{{\mathbb{P}}}(2){\raisebox{-0.1pt}{$|_{C}$}}$ is invertible (and ample) and ${\mathcal{L}}={\mathcal{O}}_{{\mathbb{P}}}(1){\raisebox{-0.1pt}{$|_{C}$}}^{[{1}]}$ is a ggs structure on $C$ satisfying the given conditions on $h^{0}$.
5.E.1. $A(0)$
Let ${\mathbb{P}}={\mathbb{P}}(2,2,3,3)$ be the weighted projective space with variables and degrees $\begin{pmatrix}y_{1}&y_{2}&z_{1}&z_{2}\\ 2&2&3&3\end{pmatrix}$ and let $C\subset{\mathbb{P}}(2,2,3,3)$ defined by the ideal
$I=(z_{1}^{2}-f_{6}(y_{1},y_{2}),z_{2}^{2}-g_{6}(y_{1},y_{2}))\ \ \mbox{ with }\ f_{6}\neq 0,\ g_{6}\neq 0.$
Computing the Hilbert series of ${\mathbb{C}}[y_{1},y_{2},z_{1},z_{2}]/I$ we see that $C$ is a curve of genus two. Moreover we have a $2:1$ map $\psi\colon C\to{\mathbb{P}}^{1}={\mathbb{P}}(2,2)$ given by $(y_{1},y_{2})$ and for a general point $P\in{\mathbb{P}}^{1}$, $\psi^{-1}(P)$ consists of two different smooth points of $C$. Therefore $C$ is reduced.
The set of points in $\mathbb{P}$ where $\mathcal{O}_{\mathbb{P}}(2)$ is not invertible is $\{y_{1}=y_{2}=0\}$ and by our assumptions it does not intersect $C$. Since
$\omega_{C}\cong\curly{E}xt^{2}(\omega_{\mathbb{P}},{\mathcal{O}}_{C})=\curly{E}xt^{2}({\mathcal{O}}_{\mathbb{P}}(-10),{\mathcal{O}}_{C}).$
and we have the resolution
$0\to{\mathcal{O}}_{\mathbb{P}}(-12)\to{\mathcal{O}}_{\mathbb{P}}(-6)^{\oplus 2}\to{\mathcal{O}}_{\mathbb{P}}\to{\mathcal{O}}_{C}\to 0$
we see that $\omega_{C}\cong\mathcal{O}_{{\mathbb{P}}}(2){\raisebox{-0.1pt}{$|_{C}$}}$ is invertible at all points of $C$, namely $C$ is Gorenstein. Finally, $h^{0}({\mathcal{L}})=h^{0}({\mathcal{O}}_{{\mathbb{P}}}(1))=0$.
5.E.2. $A(1)$
Let ${\mathbb{P}}={\mathbb{P}}(1,2,5)$ be the weighted projective space with variables and degrees $\begin{pmatrix}x&y&z\\ 1&2&5\end{pmatrix}$, and let $C\subset{\mathbb{P}}(1,2,5)$ defined by the ideal $I=(z^{2}-f_{10}(x,y))$, with $f_{10}\neq 0$.
Arguing as in the previous case, $C$ is a reduced curve of genus two and we have $\omega_{C}\cong\mathcal{O}_{{\mathbb{P}}}(2){\raisebox{-0.1pt}{$|_{C}$}}$ is invertible and $h^{0}({\mathcal{L}})=h^{0}({\mathcal{O}}_{{\mathbb{P}}}(1))=1$.
5.E.3. $B(0)$
Let ${\mathbb{P}}={\mathbb{P}}(2,2,3,3,4,5,5,6)$ be the weighted projective space with variables and degrees
$\begin{pmatrix}y_{1}&y_{2}&z_{1}&z_{2}&v&t_{1}&t_{2}&u\\ 2&2&3&3&4&5&5&6\end{pmatrix}$
and let $C$ be defined by the ideal
$I=\operatorname{rk}\begin{pmatrix}y_{1}&0&z_{1}&t_{1}\\ 0&y_{2}&z_{2}&t_{2}\\ z_{1}&z_{2}&v&u\\ t_{1}&t_{2}&u&g_{8}\end{pmatrix}\leq 1\ \ \mbox{ where }\ g_{8}=\alpha v^{2}+vh_{4}(y_{1},y_{2})+k_{8}(y_{1},y_{2})\ \mbox{ with }\alpha\neq 0$
We can write $C=C_{1}\cup C_{2}$ where $C_{i}$ is the restriction of $C$ to $y_{i}=0$.
Consider $C_{1}$: at points where $y_{2}\neq 0$ its ideal is generated by $t_{1},z_{1}$ and by the 2-by-2 minors of the matrix $\begin{pmatrix}y_{2}&z_{2}&t_{2}\\ z_{2}&v&u\\ t_{2}&u&g_{8}\end{pmatrix}$.
Computing the resolution of this ideal (e.g., taking the Eagon–Northcott complex or using MAGMA or Macaulay2), one sees that our condition on $g_{8}$ implies that it is a prime ideal, and computing the Hilbert series we see that it defines an irreducible curve $D_{1}$ with $p_{a}(D_{1})=1$. In particular $D_{1}$ is Gorenstein by [Cat82, Lemma 1.19]. An analogous computation for $C_{2}$ shows that $C$ contains $D_{1}\cup D_{2}$, where $D_{1}$ and $D_{2}$ are irreducible Gorenstein curves with arithmetic genus 1 and $C=D_{1}\cup D_{2}$ except possibly at the intersection of $C$ with $\{y_{1}=y_{2}=0\}$. This intersection consists only of the point $P:=[0,0,0,0,1,0,0,\xi]$, where $\alpha\xi^{3}=1$, and an explicit computation shows that $C$ has a node at $P$. Hence $C$ is reduced at all points, $C_{i}=D_{i}$, $i=1,2$, and $C=C_{1}\cup C_{2}$. Summing up, we have shown that $C$ is a reduced Gorenstein curve of genus two of type $B$. Hence $\omega_{C}$ is a line bundle.
Finally, we have $\omega_{C}{\raisebox{-0.1pt}{$|_{C_{i}}$}}\cong\mathcal{O}_{C_{i}}(P)\cong\mathcal{O}_{{\mathbb{P}}}(2){\raisebox{-0.1pt}{$|_{C_{i}}$}}$ (this can be seen for instance considering the hyperplane $(y_{j}=0)|_{C_{i}}$ for $i\neq j$).
Note that $C$ is contained in the locus of $\mathbb{P}$ where $\mathcal{O}_{{\mathbb{P}}}(2)$ is invertible, hence $\mathcal{O}_{{\mathbb{P}}}(2){\raisebox{-0.1pt}{$|_{C}$}}$ is a line bundle. Since the two components $C_{1}$ and $C_{2}$ of $C$ meet transversally only at one point, the natural map $\operatorname{Pic}(C)\to\operatorname{Pic}(C_{1})\times\operatorname{Pic}(C_{2})$ is an isomorphism, and therefore $\omega_{C}\cong\mathcal{O}_{{\mathbb{P}}}(2)$. [^1] Alternatively, one can use Magma or Macaulay2 to show that $\omega_{C}=\mathcal{O}_{C}(2)$. Finally we have $h^{0}({\mathcal{L}})=h^{0}({\mathcal{O}}_{{\mathbb{P}}}(1))=0$ by construction.
5.E.4. $B(1)$
Let ${\mathbb{P}}={\mathbb{P}}(1,2,3,4,5,6)$ be the weighted projective space with variables and degrees
$\begin{pmatrix}x&y&w&v&z&u\\ 1&2&3&4&5&6\end{pmatrix},$
and let $C\subset{\mathbb{P}}(1,2,3,4,5,6)$ be defined by the ideal
$\operatorname{rk}\begin{pmatrix}0&y&w&z\\ x&w&v&u\end{pmatrix}\leq 1,\begin{array}[]{rcl}z^{2}&=&yg_{8}(y,v)\\ zu&=&wg_{8}(y,v)\\ u^{2}&=&vg_{8}(y,v)+x^{4}h_{8}(x,v)\end{array}$
where $g_{8}=\alpha v^{2}+vh_{4}(y)+k_{8}(y)$ with $\alpha\neq 0$.
Arguing as in the previous case we define $C_{1}$, resp. $C_{2}$, as the restriction of $C$ to $x=0$, resp. $y=0$. Computations similar to the previous ones show that the $C_{i}$ are irreducible Gorenstein curves with $p_{a}(C_{i})=1$, they meet transversely only at the point $P:=(0:0:0:0:1:0:0:\xi)$, where $\alpha\xi^{3}=1$, and that $C$ is reduced and therefore equal to $C_{1}\cup C_{2}$. So $C$ is a Gorenstein genus 2 curve of type $B$. Finally one checks that $\omega_{C}$ is isomorphic to $\mathcal{O}_{\mathbb{P}}(2){\raisebox{-0.1pt}{$|_{C}$}}$ and $h^{0}({\mathcal{L}})=h^{0}({\mathcal{O}}_{{\mathbb{P}}}(1))=1$ by the same arguments as in case $B(0)$.
5.E.5. $B(2)$
Let ${\mathbb{P}}={\mathbb{P}}(1,1,4,6)$ be the weighted projective space with variables and degrees
$\begin{pmatrix}x_{1}&x_{2}&v&u\\ 1&1&4&6\end{pmatrix},$
and let $C\subset{\mathbb{P}}(1,1,4,6)$ defined by the ideal
[…]
One argues exactly as in cases $B(0)$ and $B(1)$ and proves that $C$ is a reduced Gorenstein curve of genus 2 that can be written $C=C_{1}\cup C_{2}$, where $C_{i}=C\cap\{x_{i}=0\}$, $i=1,2$. The curves $C_{i}$ are irreducible Gorenstein with $p_{a}(C_{i})=1$ and meet transversely at exactly one point. The line bundle $\omega_{C}$ is the restriction of $\mathcal{O}_{\mathbb{P}}(2)$ and $h^{0}({\mathcal{L}})=h^{0}({\mathcal{O}}_{{\mathbb{P}}}(1))=2$.

### v2

@@
 Recall that a spin structure (or theta-characteristic) on a smooth curve $C$ is a line bundle $\mathcal{L}$} such that there is an isomorphism $\mu\colon{\mathcal{L}}\otimes{\mathcal{L}}\to\omega_{C}$.
-The interest in such structures originated in the classical study of theta functions in the 19th century and they were taken up again in 1970’s in the context of the study of moduli spaces of curves (see [Far12]).
-They also occur in the study of algebraic surfaces: if $X$ is a smooth projective surface and $C\in|K_{X}|$ is a smooth curve in the canonical linear system, then ${\mathcal{L}}=K_{X}|_{C}$ is a theta-characteristic by the adjunction formula.
-Both in applications to moduli and to algebraic surfaces it soon became necessary to consider degenerations of the above description, with both the curve and the line bundle possibly acquiring singularities.
-In the context of moduli of curves this problem has seen some attention [Cor89, Jar98, Lud10].
-However, the application to algebraic surfaces, that sparked our interest in these structures and will be explained in a subsequent paper, requires a different approach than what was available in the literature.
-More precisely, the curves $C$ we are led to consider are Gorenstein curves $C$ of genus two with ample canonical bundle $\omega_{C}$.
-Such curves, if reduced, can be classified into two types, Type A and Type B, according to the image of the bicanonical map (Proposition 3.3).
-A generalised Gorenstein spin (ggs) structure on $C$ is given by a torsion-free sheaf ${\mathcal{L}}$ everywhere of rank one such that $\chi({\mathcal{L}})=0$ and a multiplication map $\mu\colon{\mathcal{L}}\operatorname{\otimes}{\mathcal{L}}\to\omega_{C}$ which is an isomorphism at the generic points.
-If ${\mathcal{L}}$ is in fact a line bundle, what we call a regular ggs structure, then this is the direct generalisation of the original notion.
-We will see in Section 4 that otherwise the ggs structures on $C$ are governed by regular ggs structures on partial normalisations of $C$ in many cases.
-This leads to a pretty complete understanding of the possible ggs structures on the reduced curves of interest.
-The non-reduced cases, which presents more complications, is not addressed here.
-For stable curves our definition of spin structure coincides with the definition given in [Jar98] but, in view of the application to stable surfaces, we need to include both curves with cusps and curves with elliptic tails in our analysis.
+The interest in such structures originated in the classical study of theta functions in the 19th century and they were taken up again in 1970’s in the context of the study of moduli spaces of curves (see [Far12] for a recent survey).
+They also occur in the study of algebraic surfaces: if $X$ is a smooth projective surface and $C\in|K_{X}|$ is a smooth curve in the canonical linear system, then ${\mathcal{L}}=K_{X}|_{C}$ is a theta-characteristic by the adjunction formula and the section ring $R(C,{\mathcal{L}})$ is often a stepping stone to compute the canonical ring of $X$ by Reid’s hyperplane section principle.
+In our study of the moduli space of stable surfaces, the need arose to consider a more general notion than what was available in the literature and which we hope to be of independent interest.
+We define a generalised Gorenstein spin (ggs) structure on a Gorenstein curve $C$ as a torsion-free sheaf $\mathcal{L}$} of rank one at each generic point.
+such that $\chi({\mathcal{L}})=0$ together with a multiplication map $\mu\colon{\mathcal{L}}\operatorname{\otimes}{\mathcal{L}}\to\omega_{C}$ which is an isomorphism at the generic points.
+Remark 1.1
+Let us sketch our motivation for this definition.
+Let $X$ be a stable surface $K_{X}^{2}=1$ and $h^{0}(K_{X})=2$ such that $2K_{X}$ is Cartier.
+Then any canonical curve $C\in|K_{X}|$ is a Gorenstein curve of arithmetic genus two and with ample canonical bundle by adjuction.
+However, $\omega_{X}\operatorname{\otimes}{\mathcal{O}}_{C}$ might have torsion at the non-Gorenstein points of $X$.
+So we don’t get a theta characteristic in the classical sense, but dividing out the torsion results in a ggs structure on $C$.
+The results of this paper are then an important puzzle piece in the complete classification of such surfaces $X$ in [CFPR24].
+Inspired by Mumford’s work [Mum71], Harris [Ha82] considered the case of arbitrary curves with $\mathcal{L}$} a line bundle and the multiplication map being an isomorphism.
+The Gorenstein condition is then implicit and we call this a regular ggs structure.
+We will see in Section 4 that in favourable situations, the ggs structures on a curve $C$ are governed by regular ggs structures on partial normalisations.
+In later studies of the (compactified) moduli space of spin curves, initiated by Cornalba [Cor89] and continued for example by Jarvis [Jar98] and Ludwig [Lud10], spin structures are only considered on (semi-)stable curves.
+In fact, on such curves their definition agrees with ours.
+[^1] The condition on the cokernel of the multiplication map in [Jar98, Definition 2.1.2] is automatic by Proposition 4.2, Example 2.1 and the degree condition.
+In our application to stable surfaces, Remark 1.1, we need to include both curves with cusps and curves with elliptic tails in our analysis.
 As explained in the comments after Definition 1.2 in [FS13], these two classes of curves cannot occur at the same time in a separated moduli problem for curves, so it is not clear how to deal with families in our setup.
 We have chosen not to pursue this question here.
-We then go on and analyse the half-canonical ring of such a ggs structure, that is,
+Next, we apply these results to reduced Gorenstein curves of genus two with ample canonical bundle, the class which originally motivated our research.
+In Section 3 we give a complete description of such curves, showing that we can group them in two types, Type A and Type B, distinguished by the image of the bicanonical map (Proposition 3.3) and moreover we classify all possible ggs structures.
+Such curves are hyperelliptic in the sense of [BaBo24], i.e.
+they admit a $2:1$ morphism onto a rational (not necessarily irreducible) curve (see [Ho99] for the irreducible case).
+We then go on and analyse the half-canonical ring of ggs structures on such curves, that is,
 $R(C,\{{\mathcal{L}},\omega_{C}\})=\bigoplus_{n}R_{n}$
@@
 Note that this is not a standard section ring, because the multiplication of two elements of odd degree has to be defined via the multiplication map $\mu$.
-Leaving the precise statements to Section 5 we prove the following.
-Theorem 5.2 (simplified version) —
-Let $({\mathcal{L}},\mu)$ be a ggs structure on a reduced Gorenstein curve of genus two with ample canonical bundle. Then the following cases occur:
+Leaving the precise statements to Theorem 5.2 we prove the following:
+Theorem
+Let $({\mathcal{L}},\mu)$ be a ggs structure on a reduced Gorenstein curve of genus two with ample canonical bundle.
+Then $h^{0}({\mathcal{L}})=0,1,2$ and we have a complete description of the the half-canonical ring $R(C,\{{\mathcal{L}},\omega_{C}\})$ in terms of generators and relations according to the type of the curve and to the dimension $h^{0}({\mathcal{L}})$:
 $h^{0}({\mathcal{L}})$ 0 1 2
 Type of ring $A(0)$, $B(0)$ $A(1)$, $B(1)$ $B(2)$
-In cases $A(0)$, $A(1)$, and $B(2)$ we get complete intersections, in cases $B(0)$ and $B(1)$ the rings admit a more complicated presentation in codimension six and four.
-Note that in each case we can reinterpret the ring as giving a natural embedding of $C$ into a weighted projective space ${\mathbb{P}}$, such that $\omega_{C}={\mathcal{O}}_{\mathbb{P}}(2)|_{C}$ and ${\mathcal{L}}={\mathcal{O}}_{\mathbb{P}}(1)|_{C}/\text{torsion}$.
-We hope these considerations to be of independent interest, a direct application to the moduli space of stable surfaces will be given in a subsequent paper.
+In cases $A(0)$, $A(1)$, and $B(2)$ we get complete intersections, in cases $B(0)$ and $B(1)$ the rings admit a more complicated presentation in codimension six and four (see Table 2 for a complete description).
+Note that in each case we can interpret the ring as giving a natural embedding of $C$ into a weighted projective space $\mathbb{P}$}, such that $\omega_{C}={\mathcal{O}}_{\mathbb{P}}(2)|_{C}$ and ${\mathcal{L}}={\mathcal{O}}_{\mathbb{P}}(1)|_{C}/\text{torsion}$.
 Acknowledgements
@@
 are partially supported by the project PRIN 2022BTA242 “Geometry of algebraic structures: Moduli, Invariants, Deformations” of Italian MUR and members of GNSAGA of INDAM.
+We thank the anonymous referee for some improvements of the presentation and interesting remarks.
 2. Notation and preliminary results
@@
 The sheaf $\mathcal{F}$} is reflexive if $\mathcal{F}$$\cong{\mathcal{F}}^{\vee\vee}$.
-A curve $C$ is a Cohen-Macaulay projective scheme of pure dimension 1.
+A variety is a reduced (possibly reducible) Cohen–Macaulay scheme.
+2.A.1. Curves and partial normalisation
+A curve $C$ is a Cohen–Macaulay projective scheme of pure dimension 1.
 It has a dualising sheaf $\omega_{C}$ and we denote by $p_{a}(C)$ the arithmetic genus of $C$, that is, $p_{a}(C)=1-\chi(\mathcal{O}_{C})$.
@@
 We say that $p\in C$ is a singular point of type $A_{m}$ if the completion of the local ring $\mathcal{O}_{C,p}$ is isomorphic to the ring ${\mathbb{C}}\llbracket x,y\rrbracket/(y^{2}-x^{m+1})$.
-A variety is a reduced (possibly reducible) Cohen–Macaulay scheme.
-2.B. Rank one torsion-free sheaves on Gorenstein curves
-A torsion-free sheaf on a curve $C$ is a coherent sheaf ${\mathcal{F}}$ such that $10pt({\mathcal{F}}_{x})=$ 1 for every closed point $x\in C$; equivalently ${\mathcal{F}}$ has no subsheaves supported on a finite set of points, i.e., ${\mathcal{F}}$ satisfies Serre condition $S_{1}$.
+Let $C$ be a reduced curve and $f\colon\widetilde{C}\to C$ a finite birational morphism. Then the normalisation of $C$ factors over $f$, which is therefore often called a partial normalisation.
+Note that a partial normalisation might not preserve the Gorenstein property in general: three general lines through the origin in ${\mathbb{C}}^{3}$ are not Gorenstein, but map birationally to three general lines through the origin in the plane, a Gorenstein singularity.
+For $A_{m}$ singularities these kinds of difficulties do not occur.
+Example 2.1 —
+Let us describe all partial normalisations locally at a point of type $A_{m}$ where we have two cases according to the parity of $m$, compare [Yos90, 9.9, 5.11].
+$A_{2n}$:
+The completion of the local ring is isomorphic to the ring ${\mathbb{C}}\llbracket x,y\rrbracket/(y^{2}-x^{2n+1})\cong{\mathbb{C}}\llbracket t^{2},t^{2n+1}\rrbracket$ and the partial normalisations are given by the inclusions
+${\mathbb{C}}\llbracket t\rrbracket\supset{\mathbb{C}}\llbracket t^{2},t^{3}\rrbracket\supset{\mathbb{C}}\llbracket t^{2},t^{5}\rrbracket\supset{\mathbb{C}}\llbracket t^{2},t^{7}\rrbracket\supset\dots$
+$A_{2n-1}$:
+The completion of the local ring is isomorphic to ${\mathbb{C}}\llbracket x,y\rrbracket/(y^{2}-x^{2n})$ and we get a series of inclusions
+${\mathbb{C}}\llbracket x,y\rrbracket/(y-x)\times{\mathbb{C}}\llbracket x,y\rrbracket/(y+x)\hookleftarrow{\mathbb{C}}\llbracket x,y\rrbracket/(y^{2}-x^{2})\supset{\mathbb{C}}\llbracket x,xy\rrbracket\supset{\mathbb{C}}\llbracket x,x^{2}y\rrbracket\supset\dots$
+where we consider ${\mathbb{C}}\llbracket x,x^{k}y\rrbracket$ as a subring of ${\mathbb{C}}\llbracket x,y\rrbracket/(y^{2}-x^{2})$, so that with $y_{k}=x^{k}y$ we get the desired relation $y_{k}^{2}-x^{2k+2}=0$.
+Note that in both series each successive subring has codimension one in the previous ring, when considered as a $\mathbb{C}$}-vector space.
+From this description we see that any partial normalisation of an $A_{n}$ singularity is an $A_{m}$ singularity with $m<n$ of the same parity, or smooth. In particular, it is also Gorenstein.
+2.A.2. Group action
+Assume that we have a finite group $G$ acting on a projective variety $X$ and let $\mathcal{L}$} be a rank one torsion-free sheaf on $X$.
+A linearisation of $G$ is an action of $G$ on $\mathcal{L}$} that lifts the action of $G$ on $X$.
+If $G$ is cyclic, generated by $g$, then the existence of a linearisation of $G$ is equivalent to say that there exists an isomorphism $g^{*}{\mathcal{L}}\cong{\mathcal{L}}$.
+A linearisation of $\mathcal{L}$} induces a representation of $G$ on the vector space $H^{0}(X,{\mathcal{L}})$.
+Conversely, if $\mathcal{L}$} is ample, one can recover the action of $G$ by considering its action on $R(X,{\mathcal{L}})=\bigoplus_{d\in{\mathbb{N}}}H^{0}(X,{\mathcal{L}}^{\otimes d})$.
+We will apply these notions to projective varieties having an involution, which induces a decomposition of the space of sections in invariant and anti-invariants subspaces.
+2.B. Weighted projective spaces
+We recall some facts about weighted projective spaces, see [Dol, Fle00] for details.
+Let $x_{0},\dots,x_{n}$ be coordinates on ${\mathbb{C}}^{n+1}$ and let $(w_{0},\dots,w_{n})$ be positive integers with $\gcd\{w_{1},\dots,\widehat{w}_{i},\dots,w_{n}\}=1$ for all $i=1,\dots,n$.
+Then the (well-formed) weighted projective space ${\mathbb{P}}(w_{0},\dots,w_{n})$ is defined as ${\mathbb{C}}^{n+1}/\sim$, where the equivalence relation $\sim$ is given, for $\lambda\in{\mathbb{C}}^{\ast}$, by
+$(x_{0},\dots,x_{n})\sim(\lambda^{w_{0}}x_{0},\dots,\lambda^{w_{n}}x_{n}).$
+[…]
+[…]
+$0\to{\mathcal{F}}_{c}\to\dots\to{\mathcal{F}}_{0}\to{\mathcal{O}}_{X}\to 0$
+[…]
+2.C. Rank one torsion-free sheaves on Gorenstein curves
+A torsion-free sheaf on a curve $C$ is a coherent sheaf $\mathcal{F}$} such that $0pt({\mathcal{F}}_{x})=1$ for every closed point $x\in C$; equivalently $\mathcal{F}$} has no subsheaves supported on a finite set of points, i.e., $\mathcal{F}$} satisfies Serre condition $S_{1}$.
 Convention. By a rank one torsion free sheaf on a curve $C$ we mean a torsion free sheaf that has rank one at the general point of any component of $C$.
@@
 In particular, for any such sheaf $\mathcal{F}$} we have
-[…]
+${\mathcal{F}}^{[m]}=\left({\mathcal{F}}^{\operatorname{\otimes}m}\right)^{\vee\vee}=\left({\mathcal{F}}^{\operatorname{\otimes}m}\right)/\langle\text{0-dimensional torsion}\rangle.$
 Moreover, in this case, if $\mathcal{F}$} is a rank one torsion-free sheaf and $\mathcal{H}$} is an invertible sheaf then $\deg({\mathcal{F}}\otimes{\mathcal{H}})=\deg({\mathcal{F}})+\deg({\mathcal{H}})$ (cf.
@@
 Lemma 2.4 —
-Let ${\mathcal{F}}$ be a rank one torsion free sheaf on a reduced Gorenstein curve. Then:
+Let $\mathcal{F}$} be a rank one torsion free sheaf on a reduced Gorenstein curve and assume that all the partial normalisations of $C$ are Gorenstein. Then:
   1. (i)
@@
   2. (ii)
-    $\operatorname{Aut}({\mathcal{F}})=\operatorname{Aut}(\widetilde{\mathcal{F}})$. In particular, if $\widetilde{C}$ is connected then $\operatorname{Aut}({\mathcal{F}})={\mathbb{C}}^{*}$.
+[…]
 Proof.
@@
 Lemma 2.5 —
-Let $C$ be a reduced Gorenstein curve and let $A,B\subset C$ be subcurves without common components such that $C=A\cup B$. Then there is an exact sequence:
-$0\to\omega_{B}\to\omega_{C}\to\omega_{C}{\raisebox{-0.1pt}{$|_{A}$}}\to 0$
+Let $C$ be a reduced Gorenstein curve and let $A,B\subset C$ be subcurves without common components such that $C=A\cup B$.
+Then there are a natural isomorphism ${\mathcal{I}}_{A}\omega_{C}\cong\omega_{B}$ and an exact sequence:
+$0\to\omega_{B}\to\omega_{C}\to\omega_{C}{\raisebox{-0.1pt}{$|_{A}$}}\to 0.$
 Proof.
@@
 Note that an ample line bundle has positive degree on any component, so there are at most two components and we can write $C=C_{1}\cup C_{2}$ with $\deg(\omega_{C}|_{C_{i}})=1$.
-By Lemma 2.4 there is an exact sequence
-(3.4) $0\to\omega_{C_{1}}\to\omega_{C}\to\omega_{C}|_{C_{2}}\to 0$
+By Lemma 2.5 we have an isomorphism ${\mathcal{I}}_{C_{2}}\omega_{C}\cong\omega_{C_{1}}$ and there is an exact sequence
+(3.4) $0\to\omega_{C_{1}}\to\omega_{C}\to\omega_{C}|_{C_{2}}\to 0.$
 This implies that $\omega_{C_{i}}\hookrightarrow\omega_{C}|_{C_{i}}$, so by Remark 2.2 we get $\deg\omega_{C_{i}}\leq 0$ by parity and thus $C_{i}$ has arithmetic genus at most one.
@@
 $(i)$ and $(ii)$ follow immediately by Proposition 3.3.
-$(iii)$ By Proposition 3.3 the curve $C$ is embedded in a weighted projective space $\mathbb{P}$ in such a way that the regular involution of $\mathbb{P}$ that sends $u$ to $-u$ and leaves the remaining variables invariant induces $\iota$ on $C$.
+$(iii)$ By Proposition 3.3 the curve $C$ is embedded in a weighted projective space $\mathbb{P}$ in such a way that the regular involution of $\mathbb{P}$ that sends $u$ to u-u and leaves the remaining variables invariant induces $\iota$ on $C$.
 $\square$
 Remark 3.7
-Let $X$ be a variety (so in particular a Cohen–Macaulay scheme, cf.
-section 2.A) with an involution $\iota$ that acts non-trivially on every irreducible component of $X$ and let $\pi\colon X\to Y:=X/\iota$ be the quotient map.
-Then $\pi_{*}{\mathcal{O}}_{X}$ decomposes under the action of $\iota$ as a direct sum ${\mathcal{O}}_{Y}\oplus{\mathcal{F}}$, where ${\mathcal{F}}$ is a torsion-free sheaf of rank one, and $\iota$ acts on ${\mathcal{O}}_{Y}$ as the identity and on ${\mathcal{F}}$ as multiplication by $-1$.
+Let $X$ be a projective variety, in particular reduced and Cohen–Macaulay by our conventions, with an involution $\iota$ that acts non-trivially on every irreducible component of $X$ and let $\pi\colon X\to Y:=X/\iota$ be the quotient map.
+Then $\pi_{*}{\mathcal{O}}_{X}$ decomposes under the action of $\iota$ as a direct sum ${\mathcal{O}}_{Y}\oplus{\mathcal{F}}$, where $\mathcal{F}$} is a torsion-free sheaf of rank one, and $\iota$ acts on ${\mathcal{O}}_{Y}$ as the identity and on $\mathcal{F}$} as multiplication by 1-1.
 By relative duality we have $\pi_{*}\omega_{X}=\curly{H}om(\pi_{*}{\mathcal{O}}_{X},\omega_{Y})$ and the decomposition of $\pi_{*}{\mathcal{O}}_{X}$ induces a decomposition $\pi_{*}\omega_{X}=\omega_{Y}\oplus\curly{H}om({\mathcal{F}},\omega_{Y})$, where the first summand is $\iota$-invariant and the second one is anti-invariant.
-We call the corresponding linearisation of $\omega_{Y}$ “natural”.
+We call the corresponding linearisation of $\omega_{X}$ natural.
 Note that when both $X$ and $Y$ are smooth the natural linearisation corresponds to the natural action on differential forms, for which invariant $n$-forms on $X$ correspond to pull-backs of $n$-forms on $Y$; note also that if the map $\pi$ is flat, namely $\mathcal{F}$} is a line bundle, then $\omega_{X}=\varphi^{*}(\omega_{Y}\otimes{\mathcal{F}}^{-1})$ and the opposite of the natural linearisation is the one induced by the fact that $\omega_{X}$ is a pull-back from $Y$.
@@
     • (b)
-      $C$ is reducible iff $\sum_{i}(n_{i}+1)=6$ and all the $n_{i}$ are odd. The components of $C$ are smooth rational curves exchanged by $\iota$.
+      $C$ is reducible iff $\sum_{i}(n_{i}+1)=6$ and all the $n_{i}$ are odd. In this case, the components of $C$ are smooth rational curves exchanged by $\iota$.
   2. (ii)
@@
 Type A Type B
+2 0 11
+1 0 01
+$A_{2}$1 0 01
+$A_{3} 0$0 00
+$A_{4}$0 $A_{3}0$ 00
+$A_{5}0$ 00
 4. Generalised Gorenstein spin curves
@@
 In some applications ggs structures will appear naturally also on non-reduced curves but uniform results seem to be much harder to come by, so we do not address this case here.
-Let $C$ be a reduced curve and $f\colon\widetilde{C}\to C$ a finite birational morphism.
-Then the normalisation of $C$ factors over $f$, which is therefore often called a partial normalisation.
-Note that a partial normalisation might not preserve the Gorenstein property in general: three general lines through the origin in ${\mathbb{C}}^{3}$ are not Gorenstein, but map birationally to three general lines through the origin in the plane, a Gorenstein singularity.
-Example 4.2 —
-Let us describe all partial normalisations locally at a point of type $A_{m}$ where we have two cases according to the parity of $m$, compare [Yos90, 9.9, 5.11].
-$A_{2n}$:
-The completion of the local ring is isomorphic to the ring ${\mathbb{C}}\llbracket x,y\rrbracket/(y^{2}-x^{2n+1})\cong{\mathbb{C}}\llbracket t^{2},t^{2n+1}\rrbracket$ and the partial normalisations are given by the inclusions
-${\mathbb{C}}\llbracket t\rrbracket\supset{\mathbb{C}}\llbracket t^{2},t^{3}\rrbracket\supset{\mathbb{C}}\llbracket t^{2},t^{5}\rrbracket\supset{\mathbb{C}}\llbracket t^{2},t^{7}\rrbracket\supset\dots$
-$A_{2n-1}$:
-The completion of the local ring is isomorphic to ${\mathbb{C}}\llbracket x,y\rrbracket/(y^{2}-x^{2n})$ and we get a series of inclusions
-${\mathbb{C}}\llbracket x,y\rrbracket/(y-x)\times{\mathbb{C}}\llbracket x,y\rrbracket/(y+x)\hookleftarrow{\mathbb{C}}\llbracket x,y\rrbracket/(y^{2}-x^{2})\supset{\mathbb{C}}\llbracket x,xy\rrbracket\supset{\mathbb{C}}\llbracket x,x^{2}y\rrbracket\supset\dots$
-where we consider ${\mathbb{C}}\llbracket x,x^{k}y\rrbracket$ as a subring of ${\mathbb{C}}\llbracket x,y\rrbracket/(y^{2}-x^{2})$, so that with $y_{k}=x^{k}y$ we get the desired relation $y_{k}^{2}-x^{k+2}=0$.
-Note that in both series each successive subring has codimension one in the previous ring, when considered as a ${\mathbb{C}}$-vector space.
-From this description we see that any partial normalisation of an $A_{n}$ singularity is an $A_{m}$ singularity with $m<n$ of the same parity, or smooth. In particular, it is also Gorenstein.
 The following is true if every partial normalisation of the Gorenstein curve $C$ is again Gorenstein, but we state it only for the case that we need.
 Proposition 4.2 —
-Let $({\mathcal{L}},\mu)$ be a ggs structure on a reduced Gorenstein curve with only singularities of type $A_{m}$.
-Then there exists a (uniquely determined) partial normalisation $f\colon\widetilde{C}\to C$ and a regular ggs structure (~$(\widetilde{\mathcal{L}},\widetilde{\mu})$ on $\widetilde{C}$ such that
-${\mathcal{L}}=f_{*}\widetilde{\mathcal{L}}\text{ and }{\mathcal{L}}^{[{2}]}=f_{*}\widetilde{\mathcal{L}}^{\operatorname{\otimes}2}.$
+[…]
+[…]
 Proof.
@@
 Since both sheaves have rank one and $f_{*}\widetilde{\mathcal{L}}$ is torsion-free, the induced map ${\mathcal{L}}^{[{2}]}\to f_{*}\widetilde{\mathcal{L}}^{\operatorname{\otimes}2}$ is an isomorphism.
-It remains to show that (~$(\widetilde{\mathcal{L}},\widetilde{\mu})$ is a regular ggs structure on $\widetilde{C}$.
-We have seen in Example 4.2 that $\widetilde{C}$ is again Gorenstein, so we can apply relative duality for a finite morphism of Gorenstein schemes, e.g., [Sta23, Section 0AU3], which gives
-$\operatorname{Hom}({\mathcal{L}}^{[2]},\omega_{C})=\operatorname{Hom}(f_{*}\widetilde{\mathcal{L}}^{\operatorname{\otimes}2},\omega_{C})\cong\operatorname{Hom}(\widetilde{\mathcal{L}}^{\operatorname{\otimes}2},\omega_{\widetilde{C}}),$
-and clearly $0=\chi({\mathcal{L}})=\chi(\widetilde{\mathcal{L}})$.
-Restriction to the subset where $f$ is an isomorphism shows that $\varphi\colon{\mathcal{L}}^{[{2}]}\to\omega_{C}$ is generically an isomorphism if and only if the corresponding $\widetilde{\varphi}\colon\widetilde{\mathcal{L}}^{\operatorname{\otimes}2}\to\omega_{\widetilde{C}}$ is, completing the correspondence between ggs structures.
-$\square$
+[…]
+[…]
+[…]
 With Proposition 4.2 in place, the enumeration of regular and irregular ggs structures on any given curve with type $A_{n}$ singularities is relatively simple, but the number of cases to consider grows quickly, if there are several possible partial normalisations to consider.
@@
 Proof.
-Let us write ${\mathcal{L}}=f_{*}\widetilde{\mathcal{L}}$ for a regular ggs structure (~$(\widetilde{\mathcal{L}},\widetilde{\mu})$ on a partial normalisation $f\colon\widetilde{C}\to C$.
-Recall that by Corollary 3.9 the curve $\widetilde{C}$ is also Gorenstein.
-The composition $f\circ\varphi:\widetilde{C}\to Q$ is a degree 2 finite morphism, so it induces an involution $\widetilde{\iota}$ on $\widetilde{C}$ that lifts $\iota$.
-We first observe that $\iota$ acts trivially on $\operatorname{Pic}^{0}(\widetilde{C})[2]$.
-Indeed, considering the induced action on $H^{1}({\mathcal{O}}_{\widetilde{C}})\cong H^{0}(\widetilde{C},\omega_{\widetilde{C}})$ and taking into account that $\widetilde{C}/\widetilde{\iota}$ has arithmetic genus 00 we see $\widetilde{\iota}$ acts on $H^{1}(\widetilde{C},{\mathcal{O}}_{\widetilde{C}})$ as multiplication by $-1$.
-Then the claim follows from the exponential sequence.
-It thus suffices to exhibit one regular $\widetilde{\iota}$-invariant ggs structure.
+[…]
+[…]
 By the criterion in Proposition 4.4 $(i)$ the curve $\widetilde{C}$ admits a regular ggs structure if and only if it is irreducible or $\deg\omega_{C}|_{C_{i}}$ is even on both components.
@@
 If $\widetilde{C}$ is connected but reducible, then it is the union of two smooth rational curves, meeting either in two nodes or in an $A_{3}$ point, so it has genus 1 and we can take the trivial ggs structure ${\mathcal{O}}_{\widetilde{C}}$ on $\widetilde{C}$.
-If $\widetilde{C}\neq C$ is irreducible, then it has arithmetic genus at most one and admits either a trivial or a unique ggs structure, as above.
+If $\widetilde{C}\neq C$ is irreducible, then it has arithmetic genus at most one and admits either a trivial or a unique regular ggs structure, as above.
 So we are left with the case when $C=\widetilde{C}$ is irreducible.
@@
 First observe that if $C$ is a connected Gorenstein curve of genus $g\leq 2$, then for any regular ggs structure $\mathcal{L}$} on $C$ we have $h^{0}(C,{\mathcal{L}})\leq 1$, because $\deg{\mathcal{L}}=g-1$.
-If $g=2$ and $C$ is reducible, then by Proposition 4.5 and the geometry of the curves (Proposition 3.3) any ggs structure is irregular, so we consider the corresponding regular ggs structure (~$(\widetilde{\mathcal{L}},\widetilde{\mu})$ on a partial normalisation $\widetilde{C}$.
-If $\widetilde{C}$ is connected, then $h^{0}(\widetilde{\mathcal{L}})\leq 1$ by the above.
-If $\widetilde{C}$ is disconnected, then both components have genus at most one, so $h^{0}(C,{\mathcal{L}})\leq 2$ with equality if and only if $C=C_{1}\cup C_{2}$ is of type $B$ and ${\mathcal{L}}={\mathcal{O}}_{C_{1}}\oplus{\mathcal{O}}_{C_{2}}$ is the pushforward of the unique ggs structure with a section on the components, because the unique regular ggs structure on ${\mathbb{P}}^{1}$ does not have a section.
+[…]
 $(iii)$ Follows from the fact that the points where $\mathcal{L}$} is invertible are those where $\widetilde{C}\to C$ is an isomorphism and from the fact that $\mu$ is an isomorphism for regular ggs structures.
@@
 (4.8) $0\to f_{*}({\mathcal{I}}\widetilde{\mathcal{L}})\to f_{*}\widetilde{\mathcal{L}}={\mathcal{L}}\to f_{*}(\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}})\to 0$
-The sheaf $f_{*}(\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}})$ is torsion-free and supported on $B$, so the last map in the above equation gives a surjective map ${\mathcal{L}}{\raisebox{-0.1pt}{$|_{B}$}}^{[1]}\to f_{*}(\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}})$.
-Since both sheaves are torsion-free of rank one, the map is an isomorphism.
-So we have $\chi({\mathcal{L}}{\raisebox{-0.1pt}{$|_{B}$}}^{[1]})=\chi({\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}})=\deg\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}}+1-p_{a}(\widetilde{B}})$ and $2\deg(\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}})=\deg K_{\widetilde{C}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}}$.
-Since there is an inclusion $\omega_{\widetilde{B}}\hookrightarrow\omega_{\widetilde{C}}$ (cf.
-Lemma 2.4), we get $\deg\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}}\geq p_{a}(\widetilde{B})-1\geq-1$, and if $\deg\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}}=-1$ holds then $\widetilde{B}$ is smooth rational.
+[…]
+[…]
 Assume $\deg\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}}=-1$, hence $0=\chi(\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}})=\chi({\mathcal{L}}{\raisebox{-0.1pt}{$|_{B}$}}^{[1]})$: if $B$ is not smooth rational then $p_{a}(B)=1$ and $\deg{\mathcal{L}}{\raisebox{-0.1pt}{$|_{B}$}}^{[1]}=0$.
@@
 $H^{0}(K_{C})\operatorname{\otimes}H^{0}(mK_{C})\to H^{0}((m+1)K_{C})=R_{2m+2},$
-$\ H^{0}(K_{C})\operatorname{\otimes}H^{0}\left({\mathcal{L}}\left((m-1)\right)K_{C}\right)\to H^{0}({\mathcal{L}}(mK_{C}))=R_{2m+1}$
+$H^{0}(K_{C})\operatorname{\otimes}H^{0}\left({\mathcal{L}}((m-1)K_{C})\right)\to H^{0}({\mathcal{L}}(mK_{C}))=R_{2m+1}$
 are surjective for $m\geq 3$. Hence $R$ is generated in degrees $\leq 6$.
@@
 It follows that $R$ is generated in degrees $\leq 10$.
-Since the canonical ring is generated in degrees $\leq 6$ (Proposition 3.3), we also know that there are no new generators in degree $10$square
+Since the canonical ring is generated in degrees $\leq 6$ (Proposition 3.3), we also know that there are no new generators in degree $10$.
+$\square$
 5.C. Type A
@@
 $x^{2}f_{10}(x^{2},y)+ux^{2}f_{4}(x^{2},y)=x^{2}z^{2}=\lambda^{2}y^{6}-2\lambda y^{3}u+u^{2}=\lambda^{2}y^{6}-2\lambda uy^{3}+f_{12}(x^{2},y).$
-A basis of in $R_{12}$ is $x^{12},x^{10}y,\dots,y^{6},x^{6}u,x^{4}yu,x^{2}y^{2}u,y^{3}u$, so comparing the coefficients of the monomials involving $u$ we see that $\lambda=0$, $f_{4}=0$, $xz=u$ and we have the claimed relation $z^{2}-f_{10}=0$.
+A basis of $R_{12}$ is $x^{12},x^{10}y,\dots,y^{6},x^{6}u,x^{4}yu,x^{2}y^{2}u,y^{3}u$, so comparing the coefficients of the monomials involving $u$ we see that $\lambda=0$, $f_{4}=0$, $xz=u$ and we have the claimed relation $z^{2}-f_{10}=0$.
 Comparing the Hilbert series (cf.
@@
 Comparing the Hilbert series (cf.
-Proposition 5.5) we see that there are no additional relations.
-Assume that $\alpha=\beta=\gamma=0$: then $C$ contains the point $q=[0,0,0,0,1]$ and for a general choice of $a_{6},b_{6},c_{6}$ the tangent cone to $C$ at this point has six irreducible components, so $C$ does not have an $A_{n}$ singularity at $q$, contradicting Corollary 3.9.
+Proposition 5.5) we see that there are no additional relations and we can embedd $C$ into ${\mathbb{P}}(2,2,3,3,6)$.
+Assume that $\alpha=\beta=\gamma=0$: then $C$ contains the point $q=[0,0,0,0,1]$ and for a general choice of $a_{6},b_{6},c_{6}$ the tangent cone to $C$ at this point has six irreducible components since the 3 equations give $6$ points in $\mathbb{P}(2,2,3,3)$ and the curve is the cone over these 6 points with vertex $q$, so $C$ does not have an $A_{n}$ singularity at $q$, contradicting Corollary 3.9.
 Therefore at least one of $\alpha,\beta,\gamma$ is nonzero and we can use one of the relations above to eliminate $u$.
@@
 Proof.
-By Proposition 4.3 there are a partial normalisation $f\colon\widetilde{C}\to C$ and a regular ggs structure (~$(\widetilde{\mathcal{L}},\widetilde{\mu})$ on $\widetilde{C}$ such that ${\mathcal{L}}=f_{*}\widetilde{\mathcal{L}}$.
+[…]
 $(i)$ In the proof of Proposition 4.6, $(i)$ we have seen that $\iota$ induces an involution $\widetilde{\iota}$ of $\widetilde{C}$ such that $\widetilde{\iota}^{*}\widetilde{\mathcal{L}}\cong\widetilde{\mathcal{L}}$.
@@
 Remark 3.7).
-So it is enough to show that for any given linearisation of $\widetilde{\mathcal{L}}$, the isomorphism ~~$\widetilde{\mathcal{L}}\otimes\widetilde{\mathcal{L}}\to\omega_{\widetilde{C}}$ induces on $\omega_{\widetilde{C}}$ the opposite of the natural linearisation.
+So it is enough to show that for any given linearisation of $\widetilde{\mathcal{L}}$, the isomorphism $\widetilde{\mathcal{L}}\otimes\widetilde{\mathcal{L}}\to\omega_{\widetilde{C}}$ induces on $\omega_{\widetilde{C}}$ the opposite of the natural linearisation.
 By [FFP22, Lemma 4.4] there is a trivializing open cover $\{U_{j}\}$ for $\widetilde{\mathcal{L}}$ on $\widetilde{C}$ such that the $U_{j}$ are $\widetilde{\iota}-$invariant.
-Let $\sigma_{j}$ be a local generator of ~$\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{U_{j}}$}}$ and let $q\in U_{j}$ be a point: then at least one between $\sigma_{j}+{\widetilde{\iota}}^{*}\sigma_{j}$ and $\sigma_{j}-{\widetilde{\iota}}^{*}\sigma_{j}$ does not vanish at $q$.
-So, up to refining the cover, we may assume in addition that for every $j$ the sheaf ~$\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{U_{j}}$}}$ is generated by an element $\tau_{j}$ such that ${\widetilde{\iota}}^{*}\tau_{j}=\pm\tau_{j}$.
-Via the isomorphism ~~$\widetilde{\mathcal{L}}\operatorname{\otimes}\widetilde{\mathcal{L}}\to\omega_{\widetilde{C}}$ we obtain a local trivialization of $\omega_{\widetilde{C}}$ such that for every $j$ there is a local generator $\alpha_{j}$ on $U_{j}$ that verifies ${\widetilde{\iota}}^{*}\alpha_{j}=\alpha_{j}$.
+Let $\sigma_{j}$ be a local generator of $\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{U_{j}}$}}$ and let $q\in U_{j}$ be a point: then at least one between $\sigma_{j}+{\widetilde{\iota}}^{*}\sigma_{j}$ and $\sigma_{j}-{\widetilde{\iota}}^{*}\sigma_{j}$ does not vanish at $q$.
+So, up to refining the cover, we may assume in addition that for every $j$ the sheaf $\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{U_{j}}$}}$ is generated by an element $\tau_{j}$ such that ${\widetilde{\iota}}^{*}\tau_{j}=\pm\tau_{j}$.
+Via the isomorphism $\widetilde{\mathcal{L}}\operatorname{\otimes}\widetilde{\mathcal{L}}\to\omega_{\widetilde{C}}$ we obtain a local trivialization of $\omega_{\widetilde{C}}$ such that for every $j$ there is a local generator $\alpha_{j}$ on $U_{j}$ that verifies ${\widetilde{\iota}}^{*}\alpha_{j}=\alpha_{j}$.
 In other words, $\omega_{\widetilde{C}}$ is the pull back of a line bundle $\widetilde{\mathcal{M}}$ on $\widetilde{C}/\widetilde{\iota}$ and the linearisation induced by the ggs structure coincides with the linearisation as the pullback of $\widetilde{\mathcal{M}}$.
@@
 Relative duality gives ${\varphi_{i}}_{*}\omega_{\widetilde{C}_{i}}=\omega_{Q_{i}}\oplus\omega_{Q_{i}}(a)$.
-Since $\deg\omega_{\widetilde{C}_{i}}=2\deg{\mathcal{M}}{\raisebox{-0.1pt}{$|_{Q_{i}}$}}$, we conclude that ${\mathcal{M}}{\raisebox{-0.1pt}{$|_{Q_{i}}$}}=\omega_{Q_{i}}(a)$, namely the linearisation induced on $\omega_{C_{i}}$ by the isomorphism ~~$\widetilde{\mathcal{L}}\operatorname{\otimes}\widetilde{\mathcal{L}}\to\omega_{\widetilde{C}}$ is the opposite of the natural one.
+Since $\deg\omega_{\widetilde{C}_{i}}=2\deg{\mathcal{M}}{\raisebox{-0.1pt}{$|_{Q_{i}}$}}$, we conclude that ${\mathcal{M}}{\raisebox{-0.1pt}{$|_{Q_{i}}$}}=\omega_{Q_{i}}(a)$, namely the linearisation induced on $\omega_{C_{i}}$ by the isomorphism $\widetilde{\mathcal{L}}\operatorname{\otimes}\widetilde{\mathcal{L}}\to\omega_{\widetilde{C}}$ is the opposite of the natural one.
 $\square$
@@
 Since $\mathcal{L}$} is not locally free at the intersection point $p$ of $C_{1}$ and $C_{2}$ (cf.
-Proposition 4.5), by Proposition 4.3 there is a partial normalisation $f\colon\widetilde{C}_{1}\sqcup\widetilde{C}_{2}\to C$ such that ${\mathcal{L}}=f_{*}\widetilde{\mathcal{L}}$, where ~=~$\widetilde{\mathcal{L}}=\widetilde{\mathcal{L}}_{1}\oplus\widetilde{\mathcal{L}}_{2}$ and ~$\widetilde{\mathcal{L}}_{i}$ is a line bundle supported on $\widetilde{C}_{i}$.
-Recall that $\widetilde{C}_{i}$ is Gorenstein, $p_{a}(\widetilde{C}_{i})=0$ or 1 and ~$\widetilde{\mathcal{L}}_{i}^{\otimes 2}=\omega_{\widetilde{C}_{i}}$.
+Proposition 4.4), by Proposition 4.2 there is a partial normalisation $f\colon\widetilde{C}_{1}\sqcup\widetilde{C}_{2}\to C$ such that ${\mathcal{L}}=f_{*}\widetilde{\mathcal{L}}$, where $\widetilde{\mathcal{L}}=\widetilde{\mathcal{L}}_{1}\oplus\widetilde{\mathcal{L}}_{2}$ and $\widetilde{\mathcal{L}}_{i}$ is a line bundle supported on $\widetilde{C}_{i}$.
+Recall that $\widetilde{C}_{i}$ is Gorenstein, $p_{a}(\widetilde{C}_{i})=0$ or 1 and $\widetilde{\mathcal{L}}_{i}^{\otimes 2}=\omega_{\widetilde{C}_{i}}$.
 We set ${\mathcal{L}}_{i}:=f_{*}\widetilde{\mathcal{L}}_{i}$ and $a_{i}:=h^{0}({\mathcal{L}}_{i})=h^{0}(\widetilde{\mathcal{L}}_{i})$.
@@
 Since ${\mathcal{O}}_{C_{i}}(2p_{i})=\varphi_{i}^{*}{\mathcal{O}}_{Q_{i}}(1)$, the projection formula gives
-${\varphi_{i}}_{*}{{\mathcal{L}}_{i}(np_{i})}=\begin{cases}({\varphi_{i}}_{*}{\mathcal{L}}_{i})\otimes{\mathcal{O}}_{Q_{i}}(\tfrac{n}{2})&\text{ $n$ even,}\\ ({\varphi_{i}}_{*}{\mathcal{L}}_{i}(p_{i}))\otimes{\mathcal{O}}_{Q_{i}}(\tfrac{n-1}{2})&\text{ $n$ odd.}\end{cases}$
+${\varphi_{i}}_{*}{\widetilde{\mathcal{L}}_{i}(np_{i})}=\begin{cases}({\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i})\otimes{\mathcal{O}}_{Q_{i}}(\tfrac{n}{2})&\text{ $n$ even,}\\ ({\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}(p_{i}))\otimes{\mathcal{O}}_{Q_{i}}(\tfrac{n-1}{2})&\text{ $n$ odd.}\end{cases}$
 In addition, ${\mathcal{O}}_{C_{i}}(2p_{i})$ has the pull-back linearisation, so it is enough to determine the decomposition into eigensheaves of ${\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}$ and ${\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}(p_{i})$.
-The line bundle ~$\widetilde{\mathcal{L}}_{i}(p_{i})$ has degree $p_{a}(\widetilde{C}_{i})$.
-If $p_{a}(\widetilde{C}_{i})=0$, then ~$\widetilde{\mathcal{L}}_{i}(p_{i})={\mathcal{O}}_{\widetilde{C}_{i}}$ and ${\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}(p_{i})={\mathcal{O}}_{Q_{i}}\oplus{\mathcal{O}}_{Q_{i}}(-1)$.
+The line bundle $\widetilde{\mathcal{L}}_{i}(p_{i})$ has degree $p_{a}(\widetilde{C}_{i})$.
+If $p_{a}(\widetilde{C}_{i})=0$, then $\widetilde{\mathcal{L}}_{i}(p_{i})={\mathcal{O}}_{\widetilde{C}_{i}}$ and ${\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}(p_{i})={\mathcal{O}}_{Q_{i}}\oplus{\mathcal{O}}_{Q_{i}}(-1)$.
 If $p_{a}(\widetilde{C}_{i})=1$, then $h^{0}(\widetilde{\mathcal{L}}_{i}(p_{i}))=1$ and $h^{1}(\widetilde{\mathcal{L}}_{i}(p_{i}))=0$, so that ${\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}(p_{i})={\mathcal{O}}_{Q_{i}}\oplus{\mathcal{O}}_{Q_{i}}(-1)$ also in this case.
-So for $i=1,2$ we choose the linearization of ~$\widetilde{\mathcal{L}}_{i}(p_{i})$ in such a way that ${\mathcal{O}}_{Q_{i}}$ is the invariant summand; this is the same as choosing a linearisation of ${\mathcal{L}}\otimes\omega_{C}$ as claimed in $(i)$.
+So for $i=1,2$ we choose the linearization of $\widetilde{\mathcal{L}}_{i}(p_{i})$ in such a way that ${\mathcal{O}}_{Q_{i}}$ is the invariant summand; this is the same as choosing a linearisation of ${\mathcal{L}}\otimes\omega_{C}$ as claimed in $(i)$.
 Note that choosing a linearisation of ${\mathcal{L}}\otimes\omega_{C}$ is the same as choosing a linearisation of $\mathcal{L}$}, because $\omega_{C}$ is invertible and has a chosen linearisation (cf.
@@
 Here we prove that $R(C,\{{\mathcal{L}},\omega_{C}\})$ is isomorphic to ${\mathbb{C}}[x,y,w,v,z,u]/I$ where the degrees are $\deg(x,y,w,v,z,u)=(1,2,3,4,5,6)$ and $I$ is generated by the following relations:
-2×2 minors of ($2\times 2\text{ minors of }\begin{pmatrix}0&y&w&z\\ x&w&v&u\end{pmatrix}\ \text{ and }\begin{array}[]{rcl}z^{2}&=&yg_{8}(y,v)\\ zu&=&wg_{8}(y,v)\\ u^{2}&=&vg_{8}(y,v)+x^{4}h_{8}(x,v)\end{array}$
+$2\times 2\text{ minors of }\begin{pmatrix}0&y&w&z\\ x&w&v&u\end{pmatrix}\ \text{ and }\begin{array}[]{rcl}z^{2}&=&yg_{8}(y,v)\\ zu&=&wg_{8}(y,v)\\ u^{2}&=&vg_{8}(y,v)+x^{4}h_{8}(x,v)\end{array}$
 With respect to the linearisation of Lemma 5.10 the involution $\iota$ acts on $R(C,\{{\mathcal{L}},\omega_{C}\})$ by $(x,y,w,v,z,u)\mapsto(x,y,w,v,-z,-u)$.
@@
 Since $C$ is reduced we have $x^{2}\neq 0$ so that by Proposition 3.3 and its proof we can write
-$R(C,K_{C})\cong{\mathbb{C}}[x^{2},y,v,u]/(x^{2}y,u^{2}-f_{12}(x^{2},y,v))\supset R(C,K_{C})^{+}={\mathbb{C}}[x^{2},y,v]/(x^{2}y).$
+(5.11) $R(C,K_{C})\cong{\mathbb{C}}[x^{2},y,v,u]/(x^{2}y,u^{2}-f_{12}(x^{2},y,v))\supset R(C,K_{C})^{+}={\mathbb{C}}[x^{2},y,v]/(x^{2}y).$
 Note that $xy=0$, because $x$ and $y$ are supported on different components of $C$.
@@
 Let us now look in degree $5$, where we have the decompositions
-$R_{5}^{+}\cong H^{0}({{\mathcal{O}}_{Q_{1}}})\oplus H^{0}({{\mathcal{O}}_{Q_{2}}}(1))\text{ and }R_{5}^{-}\cong H^{0}({\mathcal{O}}_{Q_{1}})\oplus H^{0}({\mathcal{O}}_{Q_{2}}(-1)).$
+[…]
 The first summand of $R_{5}^{+}$ is $\langle yw\rangle$, the second is $\langle x^{5},xv\rangle$ and we have a new anti-invariant generator $z$ of degree $5$ supported on $C_{1}$, which spans $R_{5}^{-}$.
@@
 End of the proof of Theorem 5.2: Type $B(1)$.
-We have analysed the situation up to degree $5$ already, so we start in degree $6$.
-By Proposition 3.3, $R_{6}$ is generated by $x^{2},y,v,u$ and $u$ does not vanish at the generic points of $C_{1}$ and $C_{2}$.
-Clearly, $w^{2}$ is an invariant section which is supported on $C_{1}$, so we get a relation
+We have analysed the situation up to degree $5$ already, so we start in degree $6$. By (5.11), we have
+$R_{6}^{+}=\langle x^{2},y,v\rangle\text{ and }R_{6}^{-}=\langle u\rangle$
+and $u$ is an anti-invariant section that does not vanish at the generic points of $C_{1}$ and $C_{2}$. Clearly $w^{2}$ is an invariant section which is supported on $C_{1}$, so we get a relation
 (5.13) $w^{2}=\alpha_{1}y^{3}+\alpha_{2}yv$
@@
 The two summands of $R_{7}^{-}$ are easily treated by noting that $xu\neq 0$ and $yz\neq 0$.
+In this way we have proved the independence of the generators and we can conclude by a dimension count.
 We derive the following decomposition in degree $8$ from the canonical ring:
@@
 We proceed just as we did in degree $7$.
-At the generic point of $C_{1}$, we can divide by $w$ to show that $wyv$ and $wy^{3}$ are linearly independent.
+At the generic point of $C_{1}$, we can divide by $w$ to show that wyvwyv and $wy^{3}$ are linearly independent.
 Hence the first summand of $R_{9}^{+}$ is as claimed.
@@
 $d$ $R_{d}^{+}$ $R_{d}^{-}$
-$5$ $\langle x_{1}^{5},x_{1}v\rangle\oplus\langle x_{2}^{5},x_{2}v\rangle$
-$6$ $\langle x_{1}^{6},x_{1}^{2}v\rangle\oplus\langle x_{2}^{6},x_{2}^{2}v\rangle$ $\langle u\rangle$
-$7$ $\langle x_{1}^{7},x_{1}^{3}v\rangle\oplus\langle x_{2}^{7},x_{2}^{3}v\rangle$ $\langle x_{1}u\rangle\oplus\langle x_{2}u\rangle$
-$8$ $\langle x_{1}^{8},x_{1}^{4}v\rangle\oplus\langle x_{2}^{8},x_{2}^{4}v\rangle\oplus\langle v^{2}\rangle$ $\langle x_{1}^{2}u\rangle\oplus\langle x_{2}^{2}u\rangle$
-$9$ $\langle x_{1}^{9},x_{1}^{5}v,x_{1}v^{2}\rangle\oplus\langle x_{2}^{9},x_{2}^{5}v,x_{2}v^{2}\rangle$ $\langle x_{1}^{3}u\rangle\oplus\langle x_{2}^{3}u\rangle$
+55 $\langle x_{1}^{5},x_{1}v\rangle\oplus\langle x_{2}^{5},x_{2}v\rangle$
+66 $\langle x_{1}^{6},x_{1}^{2}v\rangle\oplus\langle x_{2}^{6},x_{2}^{2}v\rangle$ $\langle u\rangle$
+77 $\langle x_{1}^{7},x_{1}^{3}v\rangle\oplus\langle x_{2}^{7},x_{2}^{3}v\rangle$ $\langle x_{1}u\rangle\oplus\langle x_{2}u\rangle$
+88 $\langle x_{1}^{8},x_{1}^{4}v\rangle\oplus\langle x_{2}^{8},x_{2}^{4}v\rangle\oplus\langle v^{2}\rangle$ $\langle x_{1}^{2}u\rangle\oplus\langle x_{2}^{2}u\rangle$
+99 $\langle x_{1}^{9},x_{1}^{5}v,x_{1}v^{2}\rangle\oplus\langle x_{2}^{9},x_{2}^{5}v,x_{2}v^{2}\rangle$ $\langle x_{1}^{3}u\rangle\oplus\langle x_{2}^{3}u\rangle$
 The direct summands group together basis elements that are supported on either one of the curves respectively both of them.
@@
 Therefore $C$ is reduced.
-The set of points in $\mathbb{P}$ where $\mathcal{O}_{\mathbb{P}}(2)$ is not invertible is $\{y_{1}=y_{2}=0\}$ and by our assumptions it does not intersect $C$. Since
+The set of points in $\mathbb{P}$ where $\mathcal{O}_{\mathbb{P}}(2)$ is not invertible is $\{y_{1}=y_{2}=0\}$ and by our assumptions it does not intersect $C$.
+By the very definition of the ideal $I$, which is generated by two polynomials of degree 6 having a relation in degree 12, we have the resolution
+$0\to{\mathcal{O}}_{\mathbb{P}}(-12)\to{\mathcal{O}}_{\mathbb{P}}(-6)^{\oplus 2}\to{\mathcal{O}}_{\mathbb{P}}\to{\mathcal{O}}_{C}\to 0.$
+By [Fle00] we have $\omega_{\mathbb{P}}={\mathcal{O}}_{{\mathbb{P}}}(-2-2-3-3)={\mathcal{O}}_{\mathbb{P}}(-10)$ and by Serre duality the subvariety $C$ of codimension two has canonical sheaf
 $\omega_{C}\cong\curly{E}xt^{2}(\omega_{\mathbb{P}},{\mathcal{O}}_{C})=\curly{E}xt^{2}({\mathcal{O}}_{\mathbb{P}}(-10),{\mathcal{O}}_{C}).$
-and we have the resolution
-$0\to{\mathcal{O}}_{\mathbb{P}}(-12)\to{\mathcal{O}}_{\mathbb{P}}(-6)^{\oplus 2}\to{\mathcal{O}}_{\mathbb{P}}\to{\mathcal{O}}_{C}\to 0$
-we see that $\omega_{C}\cong\mathcal{O}_{{\mathbb{P}}}(2){\raisebox{-0.1pt}{$|_{C}$}}$ is invertible at all points of $C$, namely $C$ is Gorenstein.
+The above resolution computes this as $\omega_{C}\cong\mathcal{O}_{{\mathbb{P}}}(2){\raisebox{-0.1pt}{$|_{C}$}}$, which is invertible at all points of $C$, namely $C$ is Gorenstein.
 Finally, $h^{0}({\mathcal{L}})=h^{0}({\mathcal{O}}_{{\mathbb{P}}}(1))=0$.
@@
 Consider $C_{1}$: at points where $y_{2}\neq 0$ its ideal is generated by $t_{1},z_{1}$ and by the 2-by-2 minors of the matrix $\begin{pmatrix}y_{2}&z_{2}&t_{2}\\ z_{2}&v&u\\ t_{2}&u&g_{8}\end{pmatrix}$.
-Computing the resolution of this ideal (e.g., taking the Eagon–Northcott complex or using MAGMA or Macaulay2), one sees that our condition on $g_{8}$ implies that it is a prime ideal, and computing the Hilbert series we see that it defines an irreducible curve $D_{1}$ with $p_{a}(D_{1})=1$.
+Computing the resolution of this ideal (e.g., taking the Eagon–Northcott complex or using MAGMA or Macaulay2, see the attached file), one sees that our condition on $g_{8}$ implies that it is a prime ideal, and computing the Hilbert series we see that it defines an irreducible curve $D_{1}$ with $p_{a}(D_{1})=1$.
 In particular $D_{1}$ is Gorenstein by [Cat82, Lemma 1.19].
 An analogous computation for $C_{2}$ shows that $C$ contains $D_{1}\cup D_{2}$, where $D_{1}$ and $D_{2}$ are irreducible Gorenstein curves with arithmetic genus 1 and $C=D_{1}\cup D_{2}$ except possibly at the intersection of $C$ with $\{y_{1}=y_{2}=0\}$.
-This intersection consists only of the point $P:=[0,0,0,0,1,0,0,\xi]$, where $\alpha\xi^{3}=1$, and an explicit computation shows that $C$ has a node at $P$.
+This intersection consists only of the point $P:=[0,0,0,0,1,0,0,\xi]$, where $\alpha\xi^{3}=1$, and an explicit computation shows that $C$ has a node at $P$ since for a generic $g_{8}$ the two curves are smooth and transverse at $P$.
 Hence $C$ is reduced at all points, $C_{i}=D_{i}$, $i=1,2$, and $C=C_{1}\cup C_{2}$.
@@
 The line bundle $\omega_{C}$ is the restriction of $\mathcal{O}_{\mathbb{P}}(2)$ and $h^{0}({\mathcal{L}})=h^{0}({\mathcal{O}}_{{\mathbb{P}}}(1))=2$.
+Data availability
+During the work on this publication, no data sets were generated, used or analyzed. Thus, there is no need for a link to a data repository.
+Declarations
+Conflict of interest. On behalf of all authors, the corresponding author states that there is no conflict of interest.